### v1

# Pap2Pat: Towards Automated Paper-to-Patent Drafting using Chunk-based Outline-guided Generation

Abstract
The patent domain is gaining attention in natural language processing research, offering practical applications in streamlining the patenting process and providing challenging benchmarks for large language models (LLMs). However, the generation of the description sections of patents, which constitute more than 90% of the patent document, has not been studied to date. We address this gap by introducing the task of outline-guided paper-to-patent generation, where an academic paper provides the technical specification of the invention and an outline conveys the desired patent structure. We present Pap2Pat, a new challenging benchmark of 1.8k patent-paper pairs with document outlines, collected using heuristics that reflect typical research lab practices. Our experiments with current open-weight LLMs and outline-guided chunk-based generation show that they can effectively use information from the paper but struggle with repetitions, likely due to the inherent repetitiveness of patent language. We release our data and code.
Data & Codehttps://github.com/boschresearch/Pap2Pat
1 Introduction
Patenting an invention is time-consuming and requires substantial knowledge of the field of the invention, as well as patent law. Therefore, patent drafting [^1] Formally, one can only draft patent applications, which become patent documents only via an office action, but for short we indistinguishably write patent. is a costly process, motivating measures to increase patent attorneys’ productivity. Some tasks in the patenting process are already technologically supported, e.g., prior art search (Shalaby and Zadrozny 2019; Stamatis 2022; Pujari et al. 2022). In contrast, patent drafting remains a predominantly manual effort. Recently, the interest of studying the capabilities of Large Language Models (LLMs) in the patent domain has increased (Shomee et al. 2024; Jiang and Goetz 2024; Casola and Lavelli 2022; Wang et al. 2024). However, prior work is limited to the generation of the sections Abstract and Claims (Hamborg et al. 2017; Lee 2020; Christofidellis et al. 2022; Lee 2023; Zuo et al. 2024; Lee 2024; Bai et al. 2024). Yet, patents also contain a patent description (comprising sections like Field of the Invention, Background, Summary, and Detailed Description) which constitutes more than 90% of the document. [^2] Measured on our dataset using the Llama-3 tokenizer: (0.7% abstract, 91.8% description, 7.5% claims) To the best of our knowledge, due to the lack of benchmark data, its automatic generation has not been studied to date despite offering both practical and scientific value. Automating this time-intensive drafting process could yield considerable cost savings. From a scientific standpoint, patent descriptions present a significant challenge for LLMs due to their technical complexity, specialized language, and substantial length.
Figure 1: Pap2Pat dataset creation. Patents and papers are matched based on heuristics. Outlines $\text{o}_{i}$ are automatically generated from patent chunks $\text{p}_{i}$ using constrained LLM decoding. Circled numbers correspond to numbers of instances.
The goal of this work is to evaluate LLMs’ proficiency in supporting the drafting of complex patent documents from a given invention specification. [^3] We do not aim to assess their capabilities to autonomously generate new inventions. In practice, the inventor typically drafts an invention report, i.e., a document describing and explaining the invention to the patent attorney as the basis for the patent drafting. In research labs, concurrent academic publishing and patenting is common practice (Murray and Stern 2005), and often the paper itself serves as an invention report. The resulting patent-paper pairs (PPPs) (Murray 2002) have been previously used to study innovation dynamics in the field of economics. While confidential free-form invention reports are not publicly available, scientific publications are. We leverage that fact to build Pap2Pat, a new benchmark of 1.8k carefully matched PPPs. Pap2Pat facilitates the study of LLMs on the complex long-text generation task of patent drafting.
Our approach for creating Pap2Pat, which contains PPPs from various domains, is illustrated in Figure 1. We design heuristics that reliably determine whether a patent and a paper describe the same invention, using metadata like author lists, dates, titles and abstracts. We apply a thorough filtering process to ensure accurate pairings and permissively licensed full texts. For LLM-supported patent drafting, we envision a practical setting in which the patent attorney, given the paper or invention report, creates an outline for the patent. This outline acts as a flexible mechanism to control the document structure and content while keeping manual effort low. This level of control also makes the approach practically usable for patent attorneys, as the effort for post-editing is also kept manageable. With Pap2Pat, we provide automatically generated outlines at varying levels of granularity.
A major challenge of the proposed task is the document length: patent descriptions in Pap2Pat are on average 18k tokens long, some exceeding 180k tokens. We propose chunk-based outline-guided generation, a novel approach to generating long patent documents chunk-by-chunk, guided by a user-defined outline and grounded in the corresponding research paper. We benchmark current open-weight LLMs on this task in both zero-shot and fine-tuning setups. Our main contributions are:
  1. (1)
    We create and release Pap2Pat, a new dataset of 1.8k PPPs including meticulously extracted full texts and several outlines for each patent.
  2. (2)
    We propose to exploit the parallelism of patents and papers to ground the generation of patent drafts in high-quality technical documents.
  3. (3)
    We propose a novel chunk-based outline-guided generation approach to patent drafting.
  4. (4)
    We conduct in-depth quantitative and qualitative analyses on Pap2Pat. While our results show encouraging performance, they also highlight limitations of current LLMs and reveal promising directions for future work.
2 Related Work
Patent-Paper Pairs (PPPs). Many research labs practice concurrent patenting and academic publishing, and this is reflected in publication and patent numbers, which are often roughly similar (Cottier, Besiroglu, and Owen 2023; Haney 2019). Murray and Stern (2005) find that almost 50% of their sampled academic papers from Nature Biotechnology have a corresponding US patent. In economics, PPPs have been used to study innovation dynamics, like whether patenting hinders the free flow of innovations (Murray and Stern 2005; Magerman, van Looy, and Debackere 2011). In that context, several approaches to finding PPPs have been proposed: Murray (2002) and Murray and Stern (2005) identify pairs manually by analyzing their full texts and citation networks. Magerman, Van Looy, and Song (2010) and Van Looy et al. (2011) explore several data mining techniques, including SVD and term-set overlaps between titles and abstracts, but they did not publish any open dataset. We build upon their idea for the document content similarity, but refine it, and add criteria for author overlaps, date ranges, competing candidates, and licenses (see Section 3.1). Gans, Murray, and Stern (2017) propose a taxonomy of PPPs that includes both to 1-to-1 matches and m-to-n matches. To ensure that papers are a solid source of information about the invention, we design our matching procedure to find only 1-to-1 matches. To the best of our knowledge, our work is the first to present a dataset of PPPs for natural language processing (NLP) research.
Patent Generation. Prior work on patent generation has focused on titles, abstracts, and claims. For instance, Christofidellis et al. (2022) train a GPT2 model to generate these parts and found that multitask learning improves performance. Lee (2023) pre-trains a GPT-J-6B architecture on patents, including the descriptions, and evaluates it on claim generation via hypothetically saved keystrokes. Zuo et al. (2024) use LLMs such as GPT-3.5-turbo and Llama-2 to generate abstracts from claims, and claims from previous claims. Jiang and Goetz (2024) provide a comprehensive review of patent generation: they identify several limitations of prior work, such as ill-posed task setups, lack of open benchmarks, and disregard for patent descriptions; all issues that we tackle in the present work.
Outline-guided Generation. Outline-guided generation is a paradigm in which LLMs use an outline to produce longer, more structured and coherent text, and which can be used to increase user control. It has been used for the generation of stories and articles, but to the best of our knowledge, it has not been applied to patent generation. The outline can be model-generated, essentially constituting a planning step. To obtain training data for this step, prior work has used extraction of keywords (Yao et al. 2019), phrases (Fang et al. 2021) and sentences (Li et al. 2023b; Sun et al. 2022; Drissi, Watkins, and Kalita 2018). Additionally, outlines could enable more sophisticated human-in-the-loop user interactions. Goldfarb-Tarrant, Feng, and Peng (2019) have shown that iteratively and interactively refining the outline yields substantial improvements in story generation. In our work, we posit that humans should provide outlines to automated patent generation methods, a modest manual effort that greatly increases user control.
Field Patent Paper Candidate 1 () Paper Candidate 2 ()
Authors Content Ge Wang, Wenxiang Cong Wenxiang Cong, Yan Xi, Bruno De Man, Ge Wang Wenxiang Cong, Yan Xi, Peter Fitzgerald, Bruno De Man, Ge Wang
$\text{sim}_{\text{author}}$ 1.0 1.0
Title Content Monochromatic CT Image Reconstruction from Current-Integrating Data via Machine Learning Monochromatic image reconstruction via machine learning Virtual Monoenergetic CT Imaging via Deep Learning
$\text{sim}_{\text{term}}$ 1.0 / 0.63 0.39 / 0.32
Abstract Content A machine-learning-based monochromatic CT image reconstruction method is described … X-ray computed tomography (CT) is a nondestructive imaging … The physical process of X-ray CT imaging is described …
$\text{sim}_{\text{term}}$ 0.71 / 0.19 0.50 / 0.23
Date 2021-08-30 2020-11-01 (✓) 2021-04-14 (✓)
Table 1: Example for the matching of a PPP. The scores correspond to the respective metrics, the term metrics are shown as min-normalized / max-normalized. Both papers have author lists that contain all the inventors, were published inside the date range, and have titles and abstracts with sufficiently high absolute term similarity to the patent. Since the term similarity scores of paper 1 are higher than those of paper 2 by the specified margin (see Distinctiveness filtering step), paper 1 is correctly matched to the patent.
3 Pap2Pat Dataset
In this section, we present the Pap2Pat dataset containing 1.8k PPPs from a variety of domains, each annotated with multiple outlines. It serves two main purposes: (1) It is an extremely challenging benchmark for LLMs because it contains very long and complex patent documents, and requires deep understanding of the technical domain as well as patent law. (2) It facilitates the development of AI-powered tools for patent drafting, where the patent attorney only performs post-editing rather than writing from scratch, potentially incurring massive cost savings. In the following, we describe the steps taken for constructing the Pap2Pat dataset: scraping and filtering PPPs, parsing the full-text documents and generating the patent outlines.
3.1 Scraping Patent-Paper Pairs
To create Pap2Pat, we start out with the USPTO dataset [^4] https://bulkdata.uspto.gov containing 6.7M patent applications from 2005 to April 2024. For each patent, we query SemOpenAlex (Färber et al. 2023) using SPARQL and retrieve papers with overlapping authors lists and publication dates. Next, we filter the results based on their titles, abstracts, other candidate matches for the same patent, and paper licenses, as elaborated below. Table 2 shows the remaining number of candidates after each filtering step. An example match is shown in Table 1. We perform a systematic manual post-hoc evaluation of the heuristics to validate their precision (see Section 3.2).
Filter Candidates
Authors + Date 930k
+ Term Overlap 100k
+ Distinctiveness 21k
+ Permissive License 1.8k
Table 2: Filter criteria and number of candidates.
Split # pairs # patent tokens # paper tokens # outline bullets
short medium long
train 1000 17808.9 $\pm$ 15132.5 7991.7 $\pm$ 4493.4 36.8 $\pm$ 29.2 73.5 $\pm$ 60.0 149.0 $\pm$ 122.0
val 242 18201.3 $\pm$ 16283.5 8018.5 $\pm$ 4089.2 37.4 $\pm$ 30.8 74.4 $\pm$ 62.9 150.6 $\pm$ 127.5
test 500 18144.9 $\pm$ 13159.5 8081.3 $\pm$ 3919.8 37.5 $\pm$ 24.7 74.9 $\pm$ 50.9 151.6 $\pm$ 103.7
nc-test 71 18415.2 $\pm$ 13995.4 9520.3 $\pm$ 4648.0 37.8 $\pm$ 22.7 76.0 $\pm$ 46.4 154.4 $\pm$ 94.8
all 1813 17977.7 $\pm$ 14724.4 8079.8 $\pm$ 4302.0 37.1 $\pm$ 28.0 74.1 $\pm$ 57.5 150.1 $\pm$ 117.0
Table 3: Dataset Statistics. Values are reported as mean $\pm$ std. Token counts correspond to the Llama-3 tokenizer.
Author Overlap. The patent and the paper of a PPP are by definition authored by overlapping sets of individuals. The overlap of author lists have therefore been identified as an effective (yet not sufficient) criteria for matching PPPs (Magerman, Van Looy, and Song 2010). The requirements for paper authorship are typically much lower than those for patent inventorship (Konski and Wu 2015). In many cases, only the main author(s) and the senior author(s) are listed as inventors. We accordingly employ an asymmetric score that only measures the fraction of inventors $i\in I$ that are also authors $a\in A$, not vice versa:
$\text{sim}_{\text{author}}$ $=\frac{|I\cap A|}{|I|}\geq 0.8$
This score’s effectiveness increases with the number of inventors, so we only consider patents with at least two inventors. Implementating $\text{sim}_{\text{author}}$ requires some form of author name disambiguation to avoid false negatives (different spellings, e.g., with, without, or with abbreviated middle name) and false positives (e.g., very common names like John Smith). There are no author identifiers shared between the patent and paper datasets, so our disambiguation uses the surface names only. To account for false negatives, we use the aliases stored in SemOpenAlex and consider an inventor to be an author if their name matches exactly with one of the aliases. False positives are marginalized by author combinations and subsequent filters: it is highly unlikely that there exist two groups of people with the same set of names working on the same topic at the same time.
Date Range. We require the paper’s publication date to be within one year before and two years after the patent application date. The former corresponds to the USPTO’s grace period, [^5] https://www.uspto.gov/web/offices/pac/mpep/s2153.html which allows inventors to file patent applications up to one year after they disclosed the invention to the public. The two-year period after the application date was selected because qualitative analyses identified it as the point of diminishing returns, beyond which the incidence of true positives notably decreases, while the rate of false positives significantly increases.
Term Overlap. We compare titles and abstracts of patents and papers using term overlap metrics. We first obtain a set of terms $T$ using removal of stopwords and punctuation666based on spaCy’s en_core_web_sm, and stemming777based on NLTK’s PorterStemmer. The score is then computed as the number of shared terms, normalized by the minimum or maximum number of terms, following Magerman, Looy, and Debackere (2015). We additionally weight each term using its IDF value across all candidate titles and abstracts. We compute these IDF weights based on the titles and abstracts of all candidate patents and papers.
$\text{sim}_{\text{term}}(s_{1},s_{2})$ $=\frac{\displaystyle\sum_{t\in T(s_{1})\cap T(s_{2})}\text{idf}(t)}{\displaystyle\text{agg}\Big{(}\sum_{t\in T(s_{1})}\text{idf}(t),\sum_{t\in T(s_{2})}\text{idf}(t)\Big{)}}$
where $s_{1}$ and $s_{2}$ are either the titles or the abstracts of the candidate pair, and $\text{agg}\in\{\min,\max\}$. Thus, we have 4 scores in total, for which we set the thresholds to be 0.15 with $\min$ normalization and 0.1 with $\max$ normalization. We choose the values based on interactive experimentation, but perform a post-hoc validation of the matching precision.
Figure 2: Our approach to chunk-based outline-guided generation. The desired section length information is used to determine chunk sizes for creating outline chunks $o_{1}$, $o_{2}$, …, $o_{n}$.
Distinctiveness. In the remaining candidate pairs, there are still many cases where one patent is matched to multiple papers or vice versa. We disambiguate these cases by comparing the term overlap metrics among these ambiguous candidate groups. We only keep a pair if 3 out of 4 term metrics are higher than those of any other candidate in the group by a margin of 0.15 and 0.1, for $\min$ and $\max$, respectively.
License. We filter the matches for licenses that allow redistribution and commercial use, i.e., CC-BY, [^8] https://creativecommons.org/licenses/by/4.0/ CC0, [^9] https://creativecommons.org/publicdomain/zero/1.0/ and public domain. We use the license information provided by SemOpenAlex and by the ArXiv API.
3.2 Manual Validation
To verify the precision of our matching pipeline, we perform a manual validation, conducted by the first author of this paper. We randomly sample 60 PPPs, read both abstracts, skim the documents, compare the figures and get an overview of the authors’ related work. We spend roughly 5 minutes per pair on average. We find that in 55/60 ($91.7\%$) pairs, the paper indeed describes the invention as one of the core contributions. In three pairs, the best match for the patent would have been a prior paper by the same authors. In two pairs, the paper would have been best matched to a related but different patent by the same inventors. This result validates the precision of our matching approach. In the five imperfect matching cases, the papers still provide meaningful training and evaluation signals, as they are still closely related to the invention and contextualized by the outline.
3.3 Document Parsing
We parse all patents and papers into a nested JSON schema where each section has a title, paragraphs and subsections field. We make considerable efforts to obtain clean data: we perform LLM-based section hierarchy reconstruction for patents, font-based section hierarchy reconstruction for paper PDFs and formula conversion for patents and paper XMLs from PubMed. We provide more details in Appendix G.
3.4 Outline Generation
One key aspect of our approach are the patent outlines, which simulate the patent attorney’s input to the system in production settings. Examples are shown in Figure 2 and Appendix A. They provide the model with the target headings, as well as bullet points summarizing the document structure and high-level content of every section.
For automatically generating the outlines for Pap2Pat, we leverage the original patents and Llama-3 70B (Dubey et al. 2024). We provide the prompt in Appendix E. To ensure that the model adheres to the output format, we use SGLang (Zheng et al. 2023) for constrained decoding. We enforce a fixed number of bullet points $n_{\text{bullets}}$ per section, where $n_{\text{bullets}}$ is proportional to the length of the text in that section ($n_{\text{chars}}$).
$n_{\text{bullets}}=\begin{cases}\text{max}(1,\lfloor n_{\text{chars}}/l\rfloor)&\text{if }n_{\text{chars}}>0\\ 0&\text{else}\end{cases}$
where $l$ is the number of characters that each bullet point summarizes on average. We create outlines in three levels of granularity: long ($l$=500, avg. 150 items), medium ($l$=1000, avg. 74 items) and short ($l$=2000, avg. 37 items), see Table 3. For each section (identified by a heading), we additionally provide the number of characters in the original patent, which can function as the desired content lengths during generation. In a practical application setting, the patent attorney would provide them alongside the outlines.
3.5 Dataset Splits and Statistics
We split our dataset randomly into train (n=1000), test (n=500) and validation (n=242). We additionally create a non-contaminated test set (nc-test) that contains all pairs with a patent published in 2024 (n=71), i.e., after the pretraining cut-off date of all evaluated open-weight LLMs. Thus, we address the concern that LLMs might have seen test data during pretraining (Ravaut et al. 2024). Table 3 shows dataset statistics across the splits, further statistics and plots are presented in Appendix H.
Dataset Split test nc-test
Metric BERTScore ROUGE-L Tokens BERTScore ROUGE-L
Outline short medium long short medium long long long long
Llama-3 8B 67.8 68.3 68.7 23.7 23.0 23.9 51.3% 68.8 22.7
Llama-3 70B 68.6 69.5 70.2 24.7 23.2 24.8 33.6% 70.5 24.1
Mixtral-8x7B 68.1 68.3 69.1 23.6 23.1 25.4 31.7% 68.8 24.3
Qwen2-72B 69.0 69.4 70.2 25.7 24.9 26.1 42.9% 70.4 25.1
Llama-3 8B SFT 69.2 69.9 70.5 26.0 27.2 26.5 162.2% 71.6 26.9
Table 4: Experimental Results. SFT stands for supervised fine-tuned. For the number of tokens and metrics on nc-test, we only show the results regarding the long outline due to space constraints.
4 Chunk-based Outline-guided Generation
In this section, we propose chunk-based outline-guided generation, a novel approach to patent generation illustrated in Figure 2. It serves as a strong baseline for comparison in future work. It is motivated by current LLMs’ limitations regarding sequence length (which needs to cover the instruction, the context, and the output) and therefore generates the patent in chunks, guided by the user-defined outline. We chunk the outline, retrieve paper context based on the outline chunk, and prompt an LLM to generate the patent text for that chunk. Finally, we concatenate the generated outputs and apply only a lightweight post-processing to remove duplicate headings at chunk boundaries.
Chunking and Token Allocation. We assign a fixed number of tokens for the instruction (including system prompt, user prompt and outlines) and equally split the remaining tokens among paper context and output patent. We then chunk the outline using the desired length of the respective patent section. The chunking procedure is designed to pack as much content as possible into each chunk, while preserving the integrity of individual sections. Therefore, we only split a section if it exceeds the token limit on its own.
Retriever. For each chunk $i$, the outline $\text{o}_{i}$ is passed to a retriever, which selects relevant paragraphs from the paper to create the paper context $\text{c}_{i}$. We use BM25 (Robertson and Zaragoza 2009) as the retrieval method and the chunk’s outline as the query. We always include the abstract of the paper because it provides a valuable overview, and successively add paragraphs in the order of their relevance ranking until the token limit is reached. The results are formatted in a Markdown document that comprises all headings and the retrieved paragraphs.
LLMs. The paper context $\text{c}_{i}$, the current outline $\text{o}_{i}$ and prior outlines $\text{o}_{j<i}$ are combined to form the prompt to the LLM (see Appendix F). Based on this prompt, the LLM generates the patent chunk $\text{p}_{i}$, where constrained decoding ensures that the model adheres strictly to the headings in the outline. To make our work reproducible, we only leverage recently published open-weights LLMs that have been reported to achieve state-of-the-art results on other generation tasks. We include a small model (Llama-3 8B (Dubey et al. 2024)) and several larger models (Mixtral-8x7B (Jiang et al. 2024), Llama-3 70B (Dubey et al. 2024) and Qwen2-72B (Yang et al. 2024)). Per default, we use a maximum sequence length of 8k tokens for all LLMs to ensure that the results are comparable and leave the study of the effect of increased maximum sequence length to future work.
Fine-tuning. We fine-tune Llama-3 8B on the training split of Pap2Pat using LoRA (Hu et al. 2022). We use the same prompt and BM25 retrieval as during inference and adopt the hyperparameters proposed by Tribes et al. (2024), see Appendix C.
5 Experiments
We perform a series of experiments using the Pap2Pat dataset to assess the performance of current open-weight LLMs on the proposed task.
5.1 Experimental Setup
To compare the generated and original patents, we use the n-gram-based metric ROUGE-L (F1) (Lin 2004), as well as the neural metric BERTScore (F1) (Zhang et al. 2020) for a balanced evaluation. For BERTScore, we use SciBERT (Beltagy, Lo, and Cohan 2019) as the base model, as it has been shown to be effective in the patent domain (Pujari, Friedrich, and Strötgen 2021). To address the context length limitations of BERT architectures, we process the input in overlapping chunks, where each chunk includes a portion of the preceding chunk to provide contextualization across chunk boundaries. We then compute the dot products between these embeddings across the whole sequence, as in the original variant. The column titled Tokens reports the ratio of the number of generated tokens and number of tokens of the original patent using the Llama-3 tokenizer.
5.2 Results
Table 4 shows the main evaluation results. The three larger models all outperform Llama-3 8B. However, all models generate patents that are much shorter than the original ones, even though the prompt includes the desired length.
BERTScore686868687070707072727272ROUGE-L2323232324242424252525252626262627272727Tokens353535354040404045454545NoPaperAbstractOnlyBM25BM25Oracle
Figure 3: Ablation of retrievers. The plot depicts the results of Qwen2-72B using the long outline.
CS (n=130) Bio (n=152)
BS RL BS RL
Llama-3 8B 67.5 22.1 69.4 25.1
Llama-3 70B 69.0 23.1 70.9 26.2
Mixtral-8x7B 67.5 23.5 70.2 26.7
Qwen2-72B 69.0 24.8 71.0 27.5
Llama-3 8B SFT 70.2 26.2 71.0 26.7
Table 5: Performance comparison between domains (CS=Computer Science, Bio=Biology). The metrics are BERTScore (BS) and ROUGE-L (RL).
Outline Granularity. With all models, the long outline achieves the highest BERTScore and ROUGE-L. We observe a monotonically increasing BERTScore with more detailed outlines. The same is not true for ROUGE-L, where the short outline consistently achieves higher scores than the medium one. Overall, the performance differences between outlines of different granularity are rather small, showcasing the robustness and flexibility of the proposed approach.
Fine-tuning. BERTScore and ROUGE-L improve substantially through fine-tuning, even beyond the scores of the larger models. It is also notable that the generated patents are more than 3 times longer than before fine-tuning. However, the increased length is at least partially due to increased proneness for generating repetitions, which we further discuss in Section 6.
Retrieval. In Figure 3, we show ablation results for the retriever. We include two baselines: NoPaper does not add any context from the paper, and AbstractOnly uses only the abstract. As an upper bound, we use BM25Oracle, where the BM25 query is the original patent text. We find that adding paper context substantially improves performance across metrics, demonstrating that associated papers provide valuable information. We observe a monotonically increasing performance across models and outline granularities. There is a small yet consistent gap between BM25 and BM25Oracle. This suggests that using the outline as a BM25 query is an effective approach, and that more elaborate retrieval methods could close that gap further.
Differences between Domains. We analyze the performance across domains and show the results in Table 5. We include the two most represented domains in the dataset: computer science and biology. All models perform slightly better on patents from biology, with the difference being more pronounced in ROUGE-L than in BERTScore.
Test Data Contamination. We use the non-contaminated test set (nc-test) to study the effect of potential pre-trained memorization on the task performance. If patents are (partially) memorized, one would expect a sudden drop in performance when evaluating on patents published after the pre-training cutoff date. We observe an equal or slightly increased BERTScore and a slightly decreased ROUGE-L, which are more likely attributed to random noise or domain shifts (see Appendix H), and do not indicate systematic issues with LLM memorization affecting our results.
6 Analysis
In this section, we provide deeper insights into the relationship between patents and papers and common error patterns.
In general, both patents and papers contain information not present in the other. The paper typically includes more experimental details and insights drawn from the experiments. The patent usually contains more information on the applications and practical benefits of the invention. We analyze the lexical overlaps between the documents and find that only 2.1% of the 4-grams are shared. This underlines the complexity of the task: the two documents describe the same invention from a different perspective using different language. Nevertheless, we find that it is common for attorneys to copy content from the paper to the patent (or vice versa). For instance, many patents and papers share a portion of the figures. But text is also commonly copied between the documents. We find that the longest common substring between paper and patent contains 27 words on average, with a maximum of 495 words. Model-generated patents and corresponding papers, on the other hand, only have a longest common substring of 15 words on average. We furthermore observe that if the paper gives the proposed method or invention a name, the model tends to use that name in the patent as well, whereas the original patents often do not.
The most prevalent error pattern in the generation process is repetition. The models sometimes get stuck in a loop where they repeat the same sentence or paragraph indefinitely. This issue is most pronounced with Llama-3 8B, in particular after fine-tuning, but also occurs with the larger models. We quantify repetitions using RR (Cettolo, Bertoldi, and Federico 2014), a metric that measures the fraction of n-grams that appear just once in a text (lower is better). The patents generated by Llama-3 8B have a much higher RR (53.2 before fine-tuning and 58.9 after fine-tuning), compared to the original patents (42.9) and larger models (Llama-3 70B: 44.1, Qwen2-72B: 40.6, Mixtral-8x7B: 42.3). We hypothesize that this issue is rooted in the inherent repetitiveness of patents. It is common that every paragraph starts with the same words for pages (e.g., ”In some embodiments, component A comprises …”). This is also reflected in the fact that patents have a much higher RR (42.9) than papers (34.5) on our test set. When generating such repetitive text, the model is more likely to get stuck in a loop of repetitions (Holtzman et al. 2020; Fu et al. 2021; Li et al. 2023a). We attempt to alleviate the issue by increasing the temperature and including a repetition penalty. We find that these approaches decrease the number of repetitions but also compromise task performance. Developing specialized methods to avoid exact-repetitions for texts that contain many near-repetitions presents an interesting direction for future research.
In our experiments, we use 8k as chunk size for all models. We also perform preliminary experiments to test whether using a larger chunk size could lead to better and more coherent patents, but find that the repetition problem overshadows these effects. For larger chunk sizes, the repetitions remain uninterrupted for longer, meaning that the content of following outline bullet points does not get addressed. We conclude that the repetition problem must be mitigated before a longer context window can achieve any benefits in automated patent drafting.
7 Conclusion
In this work, we tackle the task of patent generation. Our work is the first to study the generation of patent descriptions and to use PPPs to ground the generation process. We build the Pap2Pat benchmark dataset for training and evaluation, and propose a chunk-based approach to outline-guided paper-to-patent generation and evaluate it in both zero-shot and fine-tuning settings. We find that associated papers provide relevant information, which can be successfully injected into the LLM’s context via a simple retrieval model based on the outline. Furthermore, the approach is robust across different outline granularities and domains. We identify repetitions as the most prevalent error pattern and hypothesize that it is related to the inherent repetitiveness of patents. Promising directions for future research include addressing the repetitions, effectively leveraging models with longer context windows, and utilizing agent frameworks in the generation process.
Limitations
This study focuses on open-weight models, leaving the exploration of closed-source commercial models as a potential avenue for future research. Expanding our experimental setting to include these models could provide additional insights.
Another limitation of our work is that we have only utilized text overlap metrics, and not utilized more advanced evaluation methods, like with LLMs as judge, or via a user study. While designed to facilitate practical user interaction, the effectiveness of this approach in real-world scenarios remains to be investigated. Future studies will be necessary to fully understand the performance and user experience, and identify areas for improvement.
Ethics Statement
The development of AI-powered patent generation tools may raise ethical concerns. We emphasize that our research is intended to support and augment the work of patent professionals, rather than to automate the patenting process or facilitate malicious activities such as patent trolling.
Appendix A Patent Outline Example
1# DESCRIPTION
2
3## CROSS-REFERENCE TO RELATED APPLICATIONS
4
5- reference prior applications
6
7## BACKGROUND
8
9- limitations of current text recognition methods
10
11## SUMMARY
12
13- outline method and system for character recognition
14
15## DETAILED DESCRIPTION
16
17- introduce character recognition difficulties
18- describe lateral approach to character recognition
19- define views and bounding box
20- explain binarization and noise removal
21- describe oblique/skew detection and removal
22- outline segmentation process
23- explain lateral-view-based analysis and characteristic points selection
24- describe generation of feature vector
25- outline classification and recognition with Artificial Neural Network
26- describe training and knowledge base of Artificial Neural Network
27- summarize system and method block diagram
28
29### Handling Compound Characters
30
31- introduce compound characters
32- motivate lateral view based approach
33- discuss limitations of conventional character recognition algorithms
34- clarify scope and interpretation of patent claims
Listing 1: Example patent outline (short variant) for the pair W6364285-US20140112582. The outline corresponds to more than 5 pages. This example was randomly selected.
Appendix B Patent Structure
Figure 4: Illustration of the structure of a patent from Jiang and Goetz (2024). Note that multiple pages from the detailed description are omitted. The description includes all sections except the front matter and claims. In our experiments, we exclude sections containing only metadata, such as statements regarding funding.
Appendix C Hyperparameters
Generation Training
Parameter Value Parameter Value
max sequence length 8192 max sequence length 8192
temperature 0.6 learning rate 0.00031622
scheduler cosine
warmup ratio 0.1
epochs 3
batch size 32
lora alpha 60
lora dropout 0.05
lora r 128
Table 6: Hyperparameters during generation and training. Training parameters are adopted from Tribes et al. (2024). We run the experiments on Nvidia H100 80GB GPUs. We use a single H100 for inference and training of the 8B model and 4xH100 for the inference of the larger models using tensor parallel. We estimate the total number of GPU-hours to be 720.
Appendix D SPARQL Query
1PREFIX fabio: <http://purl.org/spar/fabio/>
2PREFIX dct: <http://purl.org/dc/terms/>
3PREFIX soa: <https://semopenalex.org/ontology/>
4PREFIX foaf: <http://xmlns.com/foaf/0.1/>
5PREFIX datacite: <http://purl.org/spar/datacite/>
6
7SELECT DISTINCT
8 ?paper ?authors ?author_names ?author_overlap ?title ?abstract ?date ?doi
[…]
13
14WHERE {
15 SELECT DISTINCT
16 ?paper
[…]
19 (SAMPLE(?author_overlap) as ?author_overlap)
20 (SAMPLE(?title) as ?title)
21 (SAMPLE(?abstract) as ?abstract)
22 (SAMPLE(?date) as ?date)
23 (SAMPLE(?doi) as ?doi)
24 ?location
25 ?url
26 ?pdf_url
27 ?license
28
29 WHERE {
30 # Information required for Matching
31 ?paper a soa:Work ;
32 dct:creator ?author ;
33 dct:title ?title ;
34 dct:abstract ?abstract ;
35 dct:date ?date ;
36 datacite:doi ?doi .
37
38 ?author foaf:name ?author_name .
39
40 # Optional Information for Downloading
41 OPTIONAL {
42 ?paper soa:hasLocation ?location .
43 OPTIONAL { ?location fabio:hasURL ?url_ . }
44 OPTIONAL { ?location soa:pdfUrl ?pdf_url_ . }
45 OPTIONAL { ?location dct:license ?license_ . }
46 BIND(COALESCE(?url_, "<EMPTY_PLACEHOLDER>") as ?url)
47 BIND(COALESCE(?pdf_url_, "<EMPTY_PLACEHOLDER>") as ?pdf_url)
48 BIND(COALESCE(?license_, "<EMPTY_PLACEHOLDER>") as ?license)
49 }
50
51 # Count number of matching authors
52 {
53 SELECT ?paper ?author_overlap
54 WHERE {
55 {
56 SELECT ?paper (COUNT(DISTINCT ?author) AS ?matching_authors)
57 WHERE {
58 ?paper dct:creator ?author .
59 ?author ?p ?name .
60 FILTER (?p IN (foaf:name, soa:alternativeName))
61 FILTER (?name IN (<AUTHOR_LIST>))
62 }
63 GROUP BY ?paper
64 ORDER BY DESC(?matching_authors)
65 LIMIT 500 # having too many results in the subquery will make query time out
66 }
67
68 ?paper dct:date ?date .
69
70 BIND (xsd:float(?matching_authors) / xsd:float(<NUM_AUTHORS>) as ?author_overlap)
71 FILTER (?author_overlap >= <AUTHOR_OVERLAP_THRESHOLD>)
72 FILTER (?date > "<DATE_EARLIEST>"^^xsd:dateTime)
73 FILTER (?date < "<DATE_LATEST>"^^xsd:dateTime)
74 }
75 }
76 }
77 GROUP BY ?paper ?location ?url ?pdf_url ?license
78 HAVING (COUNT(?author) > 1)
79}
80GROUP BY ?paper ?authors ?author_names ?author_overlap ?title ?abstract ?date ?doi
Listing 2: SPARQL query template used to retrieve papers for a given patent from SemOpenAlex. Template variables <var> are filled based on the query patent.
Appendix E Summary Generation Prompt
1<|begin_of_text|><|start_header_id|>system<|end_header_id|>
2
3You are a highly skilled patent attorney with decades of experience in drafting high-quality patent applications. You answer every question in the most concise way possible, without adding unnecessary explanations.<|eot_id|><|start_header_id|>user<|end_header_id|>
4
5### INSTRUCTION
6
7For the sections of a patent application shown below, write a bullet list that summarizes the discourse structure of the document.
8
9### OUTPUT FORMAT
10
11The output needs to be in markdown syntax. Keep the headings as they are and add bullet lists summarizing the structure of each section. Do NOT write nested lists.
12
13### GUIDELINES
14
15Here are important guidelines you need to follow:
16
17- **{n_words} words per bullet**: Every bullet point should summarize roughly {n_words} words in just a couple of words on a very high level.
18- **Structure, not content**: The bullet points should not contain all the content. You should not write a summary of the content, but a summary of the structure! For instance, you should write motivate neural networks rather than writing what the motivation for a neural network is.
19- **Start with verbal phrases**: If applicable, start the bullet points with phrases like define’, motivate’, summarize’, limitations of’, application of or embodiment’. You are not restricted to this set of phrases, just use them as inspiration. Avoid overusing the phrase describe’.
20- **Specificity**: Avoid overly generic bullet points like define method at all cost!
21- **Conciseness**: Keep every bullet point as concise as possible! Do NOT write more than 5 words per bullet!
22- **{n_bullets} bullet points in total**: You have a fixed budget of bullet points for the whole text. Make sure to write exactly {n_bullets} bullet points in total. This is with respect to all the text you are shown.
23- **Coverage**: Since you cannot write more than {n_bullets} bullet points in total, make sure you dont make the list too fine-grained in the beginning. All text must be covered! Use numbers ’(i/n)’ after the dash as progress indicators with respect to the current section.
24
25### EXAMPLE
26
27Here is an example of the output format:
28
29‘‘‘md
30# HEADING 1 (0 bullet points)
31
32## HEADING 1.1 (2 bullet points)
33
34- (1/2) introduce neural networks
35- (2/2) advantage over svm
36
37## HEADING 1.2 (3 bullet points)
38
39- (1/3) derivation of backpropagation
40- (2/3) software design of automatic differentiation
41- (3/3) example applications
42‘‘‘
43
44
45### Inputs
46
47Here is the patent application you need to summarize:
48
49
50‘‘‘md
51{context}
52‘‘‘<|eot_id|>
Listing 3: Prompt used to generate bullet point summaries with Llama-3 70B
Appendix F Patent Generation Prompt
1### ROLE
2
3You are a highly skilled patent attorney with decades of experience in drafting high-quality patent applications.
4You assist scientists in transforming their scientific discoveries into lucrative patents.
5
6### TASK DESCRIPTION
7
8Your task is to draft a patent application.
9
10### INPUTS
11
12As input, you will be provided a research paper and a patent outline, each serving a distinct purpose.
13
141. Research Paper:
15
16The research paper describes a novel invention to be patented.
17The scientist has selected the most relevant excerpts from the paper.
18Your task is to extract the invention from the paper and write a patent application for it.
19
202. Patent Outline:
21
22The patent outline summarizes the desired discourse structure of the patent document.
23It is in markdown format and contains a number of bullet points per section.
24Use this outline as a rough guidance during drafting.
25Note that the number of bullet points is also a strong indicator of the desired length! If bullet points are provided, each one corresponds to about 71 words or 1 paragraphs on average.
26You should cover all content mentioned in the outline but you are not restricted to it! Feel free to add any further information that you feel would improve the patent application.
27
283. Prior Patent Outline:
29
30Unless you are asked to generate the beginning of a patent, the user will also provide you with the outline of all prior content.
31Use it as global context where you currently stand in the process and do not repeat yourself.
32
33### GUIDELINES
34
35There are a couple of guidelines you need to follow strictly:
36
37- You might be asked to draft only parts of a patent document. Do not draft the whole patent but only those sections requested by the user.
38- Copy the headings from the outline exactly. You must include only the headings provided in the outline!
39- You must always write complete sentences and avoid keywords, bullet lists and enumerations!
40- You must use proper language and maintain a very high level of detail, as you would expect to find in a good patent!
41- The patent must act as a standalone document, therefore do not refer to the research paper in the patent!
Listing 4: System prompt used for outline-guided paper-to-patent generation
1Here are the most relevant parts of the research paper describing the invention:
2
3‘‘‘md
4# Abstract
5
6Background Heart failure patients with reduced ejection fraction (HFREF) are heterogenous, and our ability to identify patients likely to respond to therapy is limited. We present a method of identifying disease subtypes using high-dimensional clinical phenotyping and latent class analysis that may be useful in personalizing prognosis and treatment in HFREF. Methods A total of 1121 patients with nonischemic HFREF from the β-blocker Evaluation of Survival Trial were categorized according to 27 clinical features. Latent class analysis was used to generate two latent class models, LCM A and B, to identify HFREF subtypes. LCM A consisted of features associated with HF pathogenesis, whereas LCM B consisted of markers of HF progression and severity. The Seattle Heart Failure Model (SHFM) Score was also calculated for all patients. Mortality, improvement in left ventricular ejection fraction (LVEF) defined as an increase in LVEF ≥5% and a final LVEF of 35% after 12 months, and effect of bucindolol on both outcomes were compared across HFREF subtypes. Performance of models that included a combination of LCM subtypes and SHFM scores towards predicting mortality and LVEF response was estimated and subsequently validated using leave-one-out cross-validation and data from the Multicenter Oral Carvedilol Heart Failure Assessment Trial. Results A total of 6 subtypes were identified using LCM A and 5 subtypes using LCM B. Several subtypes resembled familiar clinical phenotypes. Prognosis, improvement in LVEF, and the effect of bucindolol treatment differed significantly between subtypes. Prediction improved with addition of both latent class models to SHFM for both 1-year mortality and LVEF response outcomes. Conclusions The combination of high-dimensional phenotyping and latent class analysis identifies subtypes of HFREF with implications for prognosis and response to specific therapies that may provide insight into mechanisms of disease. These subtypes may facilitate development of personalized treatment plans.
7
8
9# Introduction
10
11...
12
13We hypothesize that subtypes of nonischemic HFREF exist that may be differentiated by constellations of clinical features that reflect underlying pathophysiology. These subtypes may have variable clinical courses and responses to treatment, and identification of these subtypes may provide insight into mechanisms of HFREF and facilitate personalized prediction of outcomes and treatment response. Traditional outcomes-driven analyses are limited in the number of clinical features that can be evaluated due to the number of potential interactions between features contributing to the development and progression of HFREF. Latent class analysis is one statistical method of identifying groups of individuals within a population that share similar patterns of categorical variables such as symptoms or comorbid conditions, and it has been used in a number of medical disciplines including heart failure for exploration, characterization, and validation of diseases subtypes as well as for risk stratification and prediction of treatment response. [3]–[9] Latent class analysis has also been used to establish diagnostic standards for complex disease syndromes, and use of latent class analysis has been proposed as a method of dealing with large numbers of complex interactions and multiple comparisons in determining likelihood of response to interventions. [10]–[12] Briefly, latent class analysis hypothesizes the existence of unobserved classes within a population that explain patterns of association between variables and uses maximum-likelihood estimation to divide the population into subgroups by calculating a probability of subgroup membership for each symptom or comorbidity. An individual’s subgroup membership may therefore depend on the presence or absence of many different characteristics in a given model. When the population in question has a shared disease, the results are data-driven definitions of disease subtypes where each subtype is characterized by a distinct combination of clinical features. Many clinical variables can thereby be incorporated into an analytic model while preserving statistical power for outcomes analysis by identifying the most prevalent combinations of variables upon which to focus. We propose using complex phenotype descriptions of patients in combination with latent class analysis to identify subtypes of nonischemic HFREF that may have different prognoses and likelihoods of treatment response.
14
15...
16
17
18# Methods
19
20
21## Trial Design
22
23The design of BEST has been described previously. [14], [15] A list of all recruitment sites is found in the Appendix S1. All patients had New York Heart Association (NYHA) class III or IV HFREF (LVEF ≤35%) and were randomized in a double-blind fashion to either bucindolol or placebo. Patients were considered ischemic if they had ≥70% obstruction in a major epicardial coronary artery by angiography or evidence of prior myocardial infarction and excluded from this analysis. [16] The primary endpoint was cumulative all-cause mortality. Secondary endpoints were all-cause mortality at one year and LVEF response defined as improvement in LVEF ≥5% with a final LVEF of ≥35% as measured using multi-gated acquisition scan (MUGA). The design of MOCHA has also been described previously. [13] All patients had an LVEF ≤35%, were mostly NYHA class II or III and had stable HF symptoms for 1 month prior to enrollment. They were randomized to placebo, low (6.25 mg bid), medium (12.5 mg bid), or high-dose (25 mg bid) carvedilol. Death and LVEF improvement as measured by MUGA were secondary endpoints in the original MOCHA analysis. Mortality data was only available up to one year of follow-up in MOCHA.
Listing 5: User prompt used for outline-guided paper-to-patent generation (Part 1)
1## Identification And Definition Of Latent Classes
2
3Patients were scored according to 27 clinical features (Tables 1 and 2). Criteria were encoded and applied in a MySQL server environment (Oracle Corporation, Redwood Shores, CA). [17] Patient clinical profiles were analyzed collectively using latent class analysis [18] applied to two sets of clinical variables we designated as Latent Class Models (LCM) A and B (Tables 1 and 2). LCM A and B differed only in the clinical variables included in each model. LCM A included variables that describe a patient’s non-cardiac characteristics that can contribute to the pathogenesis of HFREF including age, gender, race, body mass index, and presence of comorbidities such as diabetes, atrial fibrillation, or valvular disease. [19]–[23] LCM B included variables that describe cardiac function, progression, and severity of HFREF including right- and left-ventricular function, hemodynamic parameters such as heart rate and blood pressure, end-organ function such as estimated creatinine clearance, and signs of venous congestion such as jugular venous distension and alanine aminotransferase levels. [24]–[33] In total, 3 variables were included in both models: body mass index, creatinine clearance, and hematocrit. All 3 variables have been implicated in the pathogenesis of HFREF and can also be markers of severity of HFREF. [34], [35] They were included in both models to illustrate that the variable implications of clinical features in different contexts may be represented using this approach. [34], [36]–[40] Two sets of related variables were also included: age of HF onset (LCM A) vs. chronologic age (LCM B) and presence of hypertension (LCM A) vs. presence of hypotension (LCM B). Age of HF onset, a static value, may be relevant to the HFREF etiology, while chronologic age may be related to HF progression. Similarly, presence of hypertension (LCM A) may be related to HF etiology while hypotension (LCM B) may be a marker of advanced HF.
4
5...
6
7
8## Association Between Latent Class Models And Outcomes
9
10...
11
12
13## Validation Of Multivariate Models
14
15...
16
17
18# Results
19
20
21## Patient Characteristics
22
23...
24
25
26## Latent Class Model A (Table 1)
27
28LCM A subtypes were characterized by distinct collections of clinical features that frequently resembled known HFREF syndromes. Subtype A1 was characterized by advanced age of onset, non-Caucasian race, male gender, HTN, mild-moderate renal insufficiency, and elevated rates of atrial fibrillation (24.5%). Subtype A2 was characterized by middle age of onset, female gender, moderate renal insufficiency, anemia, high body mass index, and very high rates of diabetes mellitus (74.6%), hypertension (95.0%), hyperlipidemia (93.8%), and hypertriglyceridemia (91.1%). Subtype A3 was characterized by middle age of onset, female gender, Caucasian race, hyperlipidemia, hypertriglyceridemia, anemia, and the presence of left bundle branch block (LBBB). Subtype A4 was characterized by young age of onset, non-Caucasian race, obesity, anemia, and lower rates of traditional cardiac risk factors such as hyperlipidemia, hypertriglyceridemia, and diabetes mellitus. Subtype A5 was characterized by advanced age of onset, Caucasian race, atrial fibrillation (86.2%), mitral valve disease (48.3%), aortic valve disease (21.8%), history of pacemaker placement (42.5%), and a significantly higher rate of prior sudden cardiac death (16.1%). This subtype had the smallest number of subjects (7.8%), whereas subtype A6 was the largest with 28.3% of subjects. Subtype A6 was characterized by middle age of onset, Caucasian race, male gender (100%), high body mass index, hypertension, hyperlipidemia, and hypertriglyceridemia with less associated diabetes mellitus (32.8%) than was seen in Subtype A2.
29
30
31## Latent Class Model B (Table 2)
32
33...
34
35
36## Association With Outcomes
37
38...
39
40
41## Differences In Treatment Effects Between Latent Classes
42
43...
44
45
46## Combined Models
47
48...
49
50
51## Model Comparisons
52
53...
54
55
56## Validation
57
58...
Listing 6: User prompt used for outline-guided paper-to-patent generation (Part 2)
1
2# Discussion
3
4Using the combination of high-dimensional clinical phenotyping and latent class analysis, we have identified a number of HFREF subtypes with distinct clinical profiles that demonstrate significant variation in prognosis as measured by all-cause mortality and response to bucindolol as measured by reduction in mortality and increased likelihood of LVEF response (Figure 4). Several of the LCM A subtypes resemble previously described nonischemic HFREF phenotypes, while LCM B subtypes model HF progression and severity. The latent class models, particularly LCM A, remained significantly associated with certain outcomes after combining them with the SHFM, suggesting that the information in the latent class models is different from the information in the SHFM Score. Taken together, these results suggest that our approach to HFREF subtype identification may be useful for identifying patients with potentially ‘reversible’ HFREF as well as those more likely to benefit from bucindolol.
5
6
7## Insight Into Mechanisms Of Disease And Treatment Response
8
9...
10
11
12## Identification Of Hfref Subtypes Using Latent Class Analysis
13
14This analysis demonstrates the potential utility of combining high-dimensional clinical phenotyping and latent class analysis for identifying relevant subtypes of HFREF. It is impossible to determine multivariate odds ratios for all of the variables included in the latent class models presented here using a traditional regression model, as the number of possible interactions (26,542,080 and 432,000,000 for LCM A and LCM B, respectively) prevents calculation using realistic sample sizes. Latent class analysis provides a quantitative mechanism of reducing the number of comparisons by aggregating individuals with similar clinical profiles. Our approach produces data-driven definitions of HFREF subtypes that integrate a large number of clinical features but are not dependent on any one feature for classification. Consequently, a feature like age may not have the same implications among all individuals. For example, subtype A4 is associated with worse outcomes than subtypes A2 or A6 despite younger age and lower burden of comorbid diseases. Clinical features may therefore be associated with a conditional probability for different outcomes depending on their context, capturing relevant interactions between comorbid conditions without direct calculation of all possible interactions. The added value of LCM A and B membership to SHFM for predicting survival despite sharing several common variables suggests that LCM A and B subtype may provide additional prognostic information to the SHFM Score. Finally, the variability in clinical outcomes observed between subtypes suggests that this approach could be useful in identifying patients with higher likelihood of HFREF reversibility in the absence of an obvious reversible etiology or conversely for identifying high risk patients for accelerated advanced HFREF therapy.
15
16
17## Implementation And Sharing
18
19...
20
21
22## Limitations
23
24...
25
26Another important limitation is the generalizability of these latent class definitions. Utilization of the coefficients derived in this analysis to determine LCM subtype for other patients assumes that the patient population is the same as the nonischemic patients enrolled in BEST. This assumption may be particularly problematic for LCM B, which includes LVEF in its definition. Like all clinical trials, the inclusion and exclusion criteria of the BEST study are a critical source of selection bias and limit the generalizability of any predictive models developed from BEST to patients that do not meet those entry criteria. [14] This is especially relevant for data-driven latent class models like those presented here, as subtype definitions are by definition dependent on the original study population, and patient subtypes not present in the derivation population might be misidentified. It must also be remembered that latent classes only represent patterns of the variables included in the models, and that those latent classes may not necessarily exist as recognizable patient types in an independent population, [6] due in part to other variables that may be important in a disease process. The utility of these models must therefore be validated further in other patient populations, and the definitions of subtypes will need to be revised over time as more diverse patient populations are incorporated.
27
28...
29
30
31## Conclusion
32
33High-dimensional phenotyping combined with latent class analysis provide a method of identifying subtypes of nonischemic HFREF patients who may have shared pathophysiology with implications for prognosis and response to bucindolol therapy. Significant reduction in all-cause mortality and increase in likelihood of LVEF response was associated with bucindolol treatment in specific groups identified using these classification methods. Identification of patients’ HFREF subtype may provide a means of personalizing clinical prognosis and estimating likelihood of responding to medical treatment.
34
35
36‘‘‘
Listing 7: User prompt used for outline-guided paper-to-patent generation (Part 3)
1Here is the outline of the desired patent application. Per bullet point, write roughly 1 paragraphs or 71 words.
2
3
4First, here is the outline of what you have already written:
5
6‘‘‘md
7# DESCRIPTION
8
9## FIELD OF THE INVENTION
10
11- define field of invention
12
13## BACKGROUND OF THE INVENTION
14
15- describe heart failure
16- limitations of current therapy
17
18## SUMMARY OF THE INVENTION
19
20- motivate invention
21- introduce HDCP and LCA
22- describe subtypes and stages of non-ischemic HF
23- correlate subtypes and stages with clinical course and β-blocker response
24- describe retrospective analysis of BEST data
25- compare results with SHFM predictions
26- assess utility of SHFM in predicting response to β-blocker
27- list exemplary β-blockers
28- describe calculated SHFM Score
29- evaluate β-blocker treatment, HF subtype, HF stage, and SHFM
30- identify 6 HF subtypes and 5 HF stages
31- associate HF subtype, HF stage, and SHFM with mortality and EF improvement
32- perform multivariate analysis
33- improve predictive performance with HF subtype and HF stage information
34- describe method for predicting response to β-blocker therapy
35- determine HF subtype and HF stage
36- use HDCP to identify HF subtypes and HF stages
37- describe method for treating non-ischemic HF patient
38- determine treatment procedure for non-ischemic HF patient
39- describe apparatus for determining HF subtype, HF stage, or combination thereof
40‘‘‘
41
42Now, continue drafting and add the following points:
43
44‘‘‘md
45# DESCRIPTION
46
47## DETAILED DESCRIPTION OF THE INVENTION
48
49- introduce non-ischemic HF patient prediction methods
50- motivate classification of patients into subtypes
51- describe clinical features influencing HF subtype and stage
52- list clinical conditions used to determine HF subtype and stage
53- outline steps to determine HF subtype
54- explain calculation of HF subtype probability
55- describe determination of HF stage
56- outline apparatuses for determining HF subtype and stage
57- describe input device for entering clinical condition information
58- explain database for storing coefficients
59- outline output device for displaying HF subtype and stage
60- discuss updating coefficients with additional data
61- describe high-throughput phenotyping method
62- outline CPU and microprocessor for calculating HF subtype and stage
63‘‘‘
64Limit your response to the sections mentioned in the summary: "DETAILED DESCRIPTION OF THE INVENTION". Remember what you have already written and do not repeat yourself.
Listing 8: User prompt used for outline-guided paper-to-patent generation (Part 4)
Appendix G Document Parsing Pipeline
In total, we use three different data sources for the full texts. For the patents, we use the USPTO bulk downloads [^10] https://bulkdata.uspto.gov/. For the papers, we use PubMed if available and PDF otherwise. The goal is to extract a clean representation of the full text into a common JSON format. In this format, every section has a title, a list of paragraphs and a list of subsections. We write parsers for the XML formats from USPTO, PubMed and the PDF parser Grobid [^11] https://github.com/kermitt2/grobid. In addition, we perform several cleaning steps:
  1. 1.
    PDF Hierarchy Reconstruction. The JSON format is hierarchical by nature, for instance to enable better chunking of patents and section-based retrieval from papers. However, Grobid does not detect section levels if the sections are not numbered. To reconstruct the levels in these cases, we implement a solution that searches for the headings in the PDF file, extracts their font properties, orders them by size, boldness and capitalization, and infers the levels from that.
  2. 2.
    Patent Hierarchy Reconstruction. In USPTO’s patent XML files, there is a level attribute associated with every heading, but we find that it is rarely correct; most headings are placed on the same level. To reconstruct the levels, we pass the list of headings to an LLM and instruct it to infer the levels based on their names (e.g., "Example 1" and "Example 2" and likely children of "EXAMPLES").
  3. 3.
    Formula Conversion. Many patents and papers include formulas that can be an important part of the document. However, in USPTO’s and PubMed’s XML formats, formulas are represented in MathML syntax, which is extremely hard to read and arguably hard to generate. To that end, we convert all MathML formulas to latex using pandoc [^12] https://github.com/jgm/pandoc.
  4. 4.
    Metadata Section Filtering. Patents usually contain a number of metadata sections in the full text, such as information regarding funding or cross-references to related patents. To filter out these sections, we collect a list of such heading names and remove a section if its heading has a Levenshtein distance less than 3 to any one of the blacklisted headings.
Appendix H Dataset Statistics
0.00.00.00.00.20.20.20.20.40.40.40.4train0.00.00.00.00.20.20.20.20.40.40.40.4val0.00.00.00.00.20.20.20.20.40.40.40.4test0.00.00.00.00.20.20.20.20.40.40.40.4nc-testBiologyComputer scienceMaterials scienceChemistryMedicinePhysicsEnvironmental scienceEngineeringothers
Figure 5: Distribution of domains across dataset splits. Domains are extracted from OpenAlex.
20012001200120012002200220022002200320032003200320042004200420042005200520052005200620062006200620072007200720072008200820082008200920092009200920102010201020102011201120112011201220122012201220132013201320132014201420142014201520152015201520162016201620162017201720172017201820182018201820192019201920192020202020202020202120212021202120222022202220222023202320232023Year0050505050100100100100150150150150200200200200FrequencyBiologyComputer scienceChemistryMedicineMaterials sciencePhysicsMathematicsEngineeringPsychologyEnvironmental scienceGeologyGeographyEconomicsBusinessPhilosophyHistoryArtPolitical science
Figure 6: Distribution of domains over time. Domains are extracted from OpenAlex.
1999200020012002200320042005200620072008200920102011201220132014201520162017201820192020202120222023Date0.00.00.00.00.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Fraction of DocumentsFraction of Candidates with PDF URL and permissive license
Figure 7: Fraction of permissive licenses over time.
400400-400- 400200200-200- 20000200200200200400400400400600600600600Days between Patent Application and Paper Publication00252525255050505075757575100100100100125125125125Number of Matches
Figure 8: Date offsets between patents and papers. Negative offset means paper was published first.

### v3

@@
-Pap2Pat: Towards Automated Paper-to-Patent Drafting using Chunk-based Outline-guided Generation
+Pap2Pat: Benchmarking Outline-Guided Long-Text Patent Generation with Patent-Paper Pairs
 Abstract
-The patent domain is gaining attention in natural language processing research, offering practical applications in streamlining the patenting process and providing challenging benchmarks for large language models (LLMs).
-However, the generation of the description sections of patents, which constitute more than 90% of the patent document, has not been studied to date.
-We address this gap by introducing the task of outline-guided paper-to-patent generation, where an academic paper provides the technical specification of the invention and an outline conveys the desired patent structure.
-We present Pap2Pat, a new challenging benchmark of 1.8k patent-paper pairs with document outlines, collected using heuristics that reflect typical research lab practices.
-Our experiments with current open-weight LLMs and outline-guided chunk-based generation show that they can effectively use information from the paper but struggle with repetitions, likely due to the inherent repetitiveness of patent language.
-We release our data and code.
-Data & Codehttps://github.com/boschresearch/Pap2Pat
+Dealing with long and highly complex technical text is a challenge for Large Language Models (LLMs), which still have to unfold their potential in supporting expensive and time-intensive processes like patent drafting.
+Within patents, the description constitutes more than 90% of the document on average.
+Yet, its automatic generation remains understudied.
+When drafting patent applications, patent attorneys typically receive invention reports (IRs), which are usually confidential, hindering research on LLM-supported patent drafting.
+Often, pre-publication research papers serve as IRs.
+We leverage this duality to build Pap2Pat, an open and realistic benchmark for patent drafting consisting of 1.8k patent-paper pairs describing the same inventions.
+To address the complex long-document patent generation task, we propose chunk-based outline-guided generation using the research paper as technical specification of the invention.
+Our extensive evaluation using Pap2Pat and a human case study show that LLMs can effectively leverage information from the paper, but still struggle to provide the necessary level of detail.
+Fine-tuning leads to more patent-style language, but also to more hallucination.
+We release our data and code at https://github.com/boschresearch/Pap2Pat.
 1 Introduction
-Patenting an invention is time-consuming and requires substantial knowledge of the field of the invention, as well as patent law.
-Therefore, patent drafting [^1] Formally, one can only draft patent applications, which become patent documents only via an office action, but for short we indistinguishably write patent.
-is a costly process, motivating measures to increase patent attorneys’ productivity.
-Some tasks in the patenting process are already technologically supported, e.g., prior art search (Shalaby and Zadrozny 2019; Stamatis 2022; Pujari et al.
-2022).
-In contrast, patent drafting remains a predominantly manual effort.
-Recently, the interest of studying the capabilities of Large Language Models (LLMs) in the patent domain has increased (Shomee et al.
-2024; Jiang and Goetz 2024; Casola and Lavelli 2022; Wang et al.
-2024).
-However, prior work is limited to the generation of the sections Abstract and Claims (Hamborg et al.
-2017; Lee 2020; Christofidellis et al.
-2022; Lee 2023; Zuo et al.
-2024; Lee 2024; Bai et al.
-2024).
-Yet, patents also contain a patent description (comprising sections like Field of the Invention, Background, Summary, and Detailed Description) which constitutes more than 90% of the document.
-[^2] Measured on our dataset using the Llama-3 tokenizer: (0.7% abstract, 91.8% description, 7.5% claims) To the best of our knowledge, due to the lack of benchmark data, its automatic generation has not been studied to date despite offering both practical and scientific value.
-Automating this time-intensive drafting process could yield considerable cost savings.
-From a scientific standpoint, patent descriptions present a significant challenge for LLMs due to their technical complexity, specialized language, and substantial length.
-Figure 1: Pap2Pat dataset creation.
-Patents and papers are matched based on heuristics.
-Outlines $\text{o}_{i}$ are automatically generated from patent chunks $\text{p}_{i}$ using constrained LLM decoding.
-Circled numbers correspond to numbers of instances.
-The goal of this work is to evaluate LLMs’ proficiency in supporting the drafting of complex patent documents from a given invention specification.
-[^3] We do not aim to assess their capabilities to autonomously generate new inventions.
-In practice, the inventor typically drafts an invention report, i.e., a document describing and explaining the invention to the patent attorney as the basis for the patent drafting.
-In research labs, concurrent academic publishing and patenting is common practice (Murray and Stern 2005), and often the paper itself serves as an invention report.
-The resulting patent-paper pairs (PPPs) (Murray 2002) have been previously used to study innovation dynamics in the field of economics.
-While confidential free-form invention reports are not publicly available, scientific publications are.
-We leverage that fact to build Pap2Pat, a new benchmark of 1.8k carefully matched PPPs.
-Pap2Pat facilitates the study of LLMs on the complex long-text generation task of patent drafting.
-Our approach for creating Pap2Pat, which contains PPPs from various domains, is illustrated in Figure 1.
-We design heuristics that reliably determine whether a patent and a paper describe the same invention, using metadata like author lists, dates, titles and abstracts.
-We apply a thorough filtering process to ensure accurate pairings and permissively licensed full texts.
-For LLM-supported patent drafting, we envision a practical setting in which the patent attorney, given the paper or invention report, creates an outline for the patent.
+Figure 1: Pap2Pat dataset creation (left) and experimental setup (right).
+Securing intellectual property is a long and costly process that requires both deep technical knowledge and expertise in patent law.
+This motivates the use of technology to boost patent attorney productivity.
+Natural language processing (NLP) already assists prior art search Shalaby and Zadrozny (2019); Stamatis (2022); Pujari et al.
+(2021) and patent landscaping (Choi et al., 2022; Pujari et al., 2022).
+Research on Large Language Models (LLMs) in the patent domain has recently gained momentum Shomee et al.
+(2024); Jiang and Goetz (2024); Casola and Lavelli (2022); Wang et al.
+(2024b), but patent drafting remains a largely manual task.
+A patent typically consists of claims, which define the invention and the legally relevant scope of protection, and a description, which provides technical details in sections like Field of the Invention, Background, Summary, and Detailed Description.
+Prior work has primarily focused on generating abstracts and claims Hamborg et al.
+(2017); Lee (2020); Christofidellis et al.
+(2022); Lee (2023); Zuo et al.
+(2024); Lee (2024); Bai et al.
+(2024b).
+The description makes up over 90% of the document,  [^1] Measured on our dataset using the Llama-3 tokenizer: 0.7% abstract, 91.8% description, 7.5% claims.
+implying that large productivity gains are expected from writing support for this section.
+Yet, generating them remains a significant challenge for LLMs due to their length, technical complexity, and specialized language Wang et al.
+(2024a, c).
+Existing work on automatic patent generation suffers from a number of shortcomings, such as ill-posed task setups, lack of open benchmarks, and disregard for patent descriptions (Jiang and Goetz, 2024).
+In this paper, we tackle all of these issues: we propose a novel setup, evaluation metrics, and outline-guided models for description generation in a realistic setting using open and human-created data.
+Inventors typically submit invention reports (IRs), which patent attorneys formalize into patent applications.
+In many research labs, it is common to use a pre-publication paper as invention report, which leads to so-called patent-paper pairs (PPPs, Murray (2002)), an unrecognized treasure for AI research.
+To facilitate the study of LLMs on patent drafting, we create Pap2Pat, a new benchmark of 1.8k carefully identified PPPs and patent outlines.
+We develop and validate a method for reliably matching patents and papers describing the same invention (see Figure 1).
+For LLM-supported patent drafting, we envision a practical setting in which the attorney, given a paper, provides an outline for the patent.
 This outline acts as a flexible mechanism to control the document structure and content while keeping manual effort low.
-This level of control also makes the approach practically usable for patent attorneys, as the effort for post-editing is also kept manageable.
-With Pap2Pat, we provide automatically generated outlines at varying levels of granularity.
-A major challenge of the proposed task is the document length: patent descriptions in Pap2Pat are on average 18k tokens long, some exceeding 180k tokens.
-We propose chunk-based outline-guided generation, a novel approach to generating long patent documents chunk-by-chunk, guided by a user-defined outline and grounded in the corresponding research paper.
-We benchmark current open-weight LLMs on this task in both zero-shot and fine-tuning setups.
+A major challenge of the proposed task is document length: patent descriptions in Pap2Pat are on average 18k tokens long, some exceeding 180k tokens.
+While current LLMs increasingly support long context windows, they struggle to generate similarly long outputs Liu et al.
+(2024); Bai et al.
+(2024a); Wu et al.
+(2025); Ye et al.
+(2025).
+As a remedy, we propose chunk-based outline-guided patent generation (COPGen), which effectively generates long patent documents in chunks.
+Evaluating generated patents poses significant challenges due to their length, their technical complexity, and the high cost of manual evaluation.
+No standard evaluation metrics are established in the literature, and prior work Wang et al.
+(2024c, a) commonly resorts to standard text similarity metrics which do not work well on long documents Lattimer et al.
+(2023); Que et al.
+(2024).
+We adapt a suite of metrics based on natural language inference (NLI) and authorship attribution to the specific case of evaluating factual correctness, coverage, and language style for long-form patent generation on Pap2Pat.
 Our main contributions are:
   1. (1)
-    We create and release Pap2Pat, a new dataset of 1.8k PPPs including meticulously extracted full texts and several outlines for each patent.
+    We create and release Pap2Pat, a new benchmark for patent drafting based on open data that closely aligns with real-world settings.
   2. (2)
-    We propose to exploit the parallelism of patents and papers to ground the generation of patent drafts in high-quality technical documents.
+    We derive a comprehensive suite of automatic evaluation metrics for evaluating generated patents.
   3. (3)
-    We propose a novel chunk-based outline-guided generation approach to patent drafting.
+    We propose a chunk-based outline-guided patent generation approach with effectively controllable output length. Our method increases coverage considerably while keeping factuality high.
   4. (4)
-    We conduct in-depth quantitative and qualitative analyses on Pap2Pat.
-While our results show encouraging performance, they also highlight limitations of current LLMs and reveal promising directions for future work.
+    We conduct extensive evaluations finding that state-of-the-art LLMs can effectively use information from the papers, but still struggle to provide the necessary level of detail, and that fine-tuning leads to much higher stylistic similarity with patents, but substantially decreased factuality.
+  5. (5)
+    Our human evaluation confirms these findings and indicates promising potential for productivity gains of patent attorneys.
 2 Related Work
 Patent-Paper Pairs (PPPs).
-Many research labs practice concurrent patenting and academic publishing, and this is reflected in publication and patent numbers, which are often roughly similar (Cottier, Besiroglu, and Owen 2023; Haney 2019).
+Many research labs practice concurrent patenting and academic publishing.
 Murray and Stern (2005) find that almost 50% of their sampled academic papers from Nature Biotechnology have a corresponding US patent.
-In economics, PPPs have been used to study innovation dynamics, like whether patenting hinders the free flow of innovations (Murray and Stern 2005; Magerman, van Looy, and Debackere 2011).
+In economics, PPPs have been used to study innovation dynamics, like whether patenting promotes or hinders the free flow of innovations Murray and Stern (2005); Magerman et al.
+(2011).
 In that context, several approaches to finding PPPs have been proposed: Murray (2002) and Murray and Stern (2005) identify pairs manually by analyzing their full texts and citation networks.
-Magerman, Van Looy, and Song (2010) and Van Looy et al.
-(2011) explore several data mining techniques, including SVD and term-set overlaps between titles and abstracts, but they did not publish any open dataset.
-We build upon their idea for the document content similarity, but refine it, and add criteria for author overlaps, date ranges, competing candidates, and licenses (see Section 3.1).
-Gans, Murray, and Stern (2017) propose a taxonomy of PPPs that includes both to 1-to-1 matches and m-to-n matches.
+Magerman et al.
+(2010) and Van Looy et al.
+(2011) explore several data mining techniques.
+We refine and extend their approach, and add criteria for author overlaps, date ranges, competing candidates, and licenses.
+Gans et al.
+(2017) propose a taxonomy of PPPs that includes both to 1-to-1 matches and m-to-n matches.
 To ensure that papers are a solid source of information about the invention, we design our matching procedure to find only 1-to-1 matches.
-To the best of our knowledge, our work is the first to present a dataset of PPPs for natural language processing (NLP) research.
+We publish, to the best of our knowledge, the first PPPs dataset for NLP research.
+Split # pairs # patent tokens # paper tokens # outline bullets short medium long train 1000 17.8k $\pm$ 15.1k 7.9k $\pm$ 4.5k 36.8 $\pm$ 29.2 73.5 $\pm$ 60.0 149.0 $\pm$ 122.0 val 242 18.2k $\pm$ 16.3k 8.0k $\pm$ 4.1k 37.4 $\pm$ 30.8 74.4 $\pm$ 62.9 150.6 $\pm$ 127.5 test 500 18.1k $\pm$ 13.2k 8.1k $\pm$ 3.9k 37.5 $\pm$ 24.7 74.9 $\pm$ 50.9 151.6 $\pm$ 103.7 nc-test 71 18.4k $\pm$ 13.9k 9.5k $\pm$ 4.6k 37.8 $\pm$ 22.7 76.0 $\pm$ 46.4 154.4 $\pm$ 094.8 all 1813 17.9k $\pm$ 14.7k 8.1k $\pm$ 4.3k 37.1 $\pm$ 28.0 74.1 $\pm$ 57.5 150.1 $\pm$ 117.0 Table 1: Pap2Pat statistics.
+Values are reported as mean $\pm$ std.
+Token counts correspond to the Llama-3 tokenizer.
 Patent Generation.
-Prior work on patent generation has focused on titles, abstracts, and claims.
-For instance, Christofidellis et al.
-(2022) train a GPT2 model to generate these parts and found that multitask learning improves performance.
-Lee (2023) pre-trains a GPT-J-6B architecture on patents, including the descriptions, and evaluates it on claim generation via hypothetically saved keystrokes.
+Most prior work on patent generation has focused on titles, abstracts, and claims.
+Christofidellis et al.
+(2022) train a multi-task GPT2 model to generate these parts.
+Lee (2023) pre-trains a GPT-J-6B architecture on entire patents and evaluates it on claim generation.
 Zuo et al.
-(2024) use LLMs such as GPT-3.5-turbo and Llama-2 to generate abstracts from claims, and claims from previous claims.
-Jiang and Goetz (2024) provide a comprehensive review of patent generation: they identify several limitations of prior work, such as ill-posed task setups, lack of open benchmarks, and disregard for patent descriptions; all issues that we tackle in the present work.
-Outline-guided Generation.
-Outline-guided generation is a paradigm in which LLMs use an outline to produce longer, more structured and coherent text, and which can be used to increase user control.
-It has been used for the generation of stories and articles, but to the best of our knowledge, it has not been applied to patent generation.
-The outline can be model-generated, essentially constituting a planning step.
-To obtain training data for this step, prior work has used extraction of keywords (Yao et al.
-2019), phrases (Fang et al.
-2021) and sentences (Li et al.
-2023b; Sun et al.
-2022; Drissi, Watkins, and Kalita 2018).
-Additionally, outlines could enable more sophisticated human-in-the-loop user interactions.
-Goldfarb-Tarrant, Feng, and Peng (2019) have shown that iteratively and interactively refining the outline yields substantial improvements in story generation.
-In our work, we posit that humans should provide outlines to automated patent generation methods, a modest manual effort that greatly increases user control.
+(2024) use GPT-3.5-turbo and Llama-2 to generate abstracts from claims, and claims from previous claims.
+Wang et al.
+(2024a) experiment with the generation of individual description paragraphs based on patent claims and drawing descriptions.
+In practice, it is not likely that claims are already finalized at the time of writing of the description section.
+In their work concurrent to ours, Wang et al.
+(2024c) leverage an agent framework to generate complete patents based on purely automatically generated invention specifications.
+They share our core principle of using a divide-and-conquer strategy for long-document generation, but evaluate only using surface-level metrics.
+Our more realistic setting relies on real-world invention specifications and provides deeper insights due to more sophisticated evaluation metrics.
+Outline-guided Generation is a paradigm in which LLMs use an outline to produce longer, more structured and coherent text.
+Outlines are either generated in a planning stage Drissi et al.
+(2018); Yao et al.
+(2019); Sun et al.
+(2022); Yang et al.
+(2022); Lee et al.
+(2024); Wang et al.
+(2024d); Shao et al.
+(2024) or provided as input Fang et al.
+(2021); Spangher et al.
+(2022); Yang et al.
+(2023); Li et al.
+(2023b).
+They are created using extraction of key words Yao et al.
+(2019), phrases Fang et al.
+(2021), or sentences Drissi et al.
+(2018); Sun et al.
+(2022); Yang et al.
+(2022); Li et al.
+(2023b), generated using LLMs Yang et al.
+(2023) or defined interactively Goldfarb-Tarrant et al.
+(2019).
+In our work, we posit that outlines should be provided by the patent attorney to satisfy the high demand for user control.
+3 Pap2Pat Benchmark
+We present the Pap2Pat dataset containing 1.8k PPPs, each annotated with three outlines for generation.
+We describe the steps taken to construct Pap2Pat, analyse the obtained corpus, and propose evaluation metrics for our new benchmark.
+3.1 Dataset Construction
+We here give an overview of the construction of Pap2Pat (for details, see Appendix A).
+Scraping PPPs.
+The patent and paper in a PPP typically do not cite each other, so we cannot rely on front-page or in-text citations Marx and Fuegi (2020, 2022) to find PPPs.
+Therefore, prior work has developed heuristics to match patents and papers based on document metadata Magerman et al.
+(2010); Van Looy et al.
+(2011).
+We use the USPTO dataset [^2] https://bulkdata.uspto.gov containing 6.7M patent applications from 2005 to April 2024.
+For each patent, we query SemOpenAlex Färber et al.
+(2023) using SPARQL and retrieve papers with overlapping authors lists and publication dates.
+We filter the results based on titles, abstracts, other candidate matches for the same patent, and paper licenses.
+Table 2 shows the remaining number of candidates after each filtering step.
+A major limiting factor for our benchmark size are the restrictive licenses of many scientific articles.
+[^3] ACL with its CC-BY license being a positive exception.
+Manual Validation.
+To verify the precision of our matching pipeline, the first author performs a manual validation.
+We randomly sample 60 PPPs, read both abstracts, skim the documents, compare the figures and get an overview of the authors’ related work.
+We spend a total of 5 hours, i.e., 5 minutes per pair on average.
+In 55/60 ($91.7\%$) pairs, the paper describes the invention as a core contribution.
+In three pairs, the best match for the patent would have been a prior paper by the same authors.
+In two pairs, the paper would have been best matched to a related but different patent by the same inventors.
+In these five imperfect matching cases, the papers still provide meaningful training and evaluation signals, as they are highly relevant to the invention and contextualized by the outline.
+Overall, the precision of our matching approach is high.
+Outline Generation.
+Outlines consist of target headings and short bullet points summarizing the document structure and high-level content of every section (see Figure 2 and Appendix E).
+For Pap2Pat, we generate them automatically from the original patents using Llama-3 70B Dubey et al.
+(2024).
+In practice, they are provided by the user.
+To ensure that the model adheres to the desired output format, we use SGLang Zheng et al.
+(2023) for constrained decoding.
+We enforce a fixed number of bullet points $n_{\text{bullets}}$ per section, where $n_{\text{bullets}}$ is proportional to the length of the text in that section ($n_{\text{chars}}$) and defined as
+$n_{\text{bullets}}=\begin{cases}\text{max}(1,\lfloor n_{\text{chars}}/l\rfloor)&\text{if }n_{\text{chars}}>0\\ 0&\text{else}\end{cases}$
+where $l$ is the number of characters that each bullet point summarizes on average.
+We create outlines with three levels of granularity: long ($l$=500, avg.
+150 items), medium ($l$=1000, avg.
+74 items) and short ($l$=2000, avg.
+37 items), see Table 1.
+The average bullet point length is 5.4 $\pm$ 2.3 words.
+For each section, we additionally provide the number of characters in the original patent to signal the desired content lengths during generation.
+Filter # pairs
+Authors + Date 930k
++ Term Overlap 100k
++ Distinctiveness 21k
++ Permissive License 1.8k
+Table 2: Filter criteria and remaining number of candidates.
+3.2 Task Description and Data Splits
+We propose the following generation task for Pap2Pat: Given a patent outline $O$ and a research paper $C$ containing a specification of the invention, models should output a patent $P$.
+The input data provides the desired output length in characters per section.
+We split our dataset randomly into train (n=1000), test (n=500) and validation (n=242), see Table 1.
+We additionally create a non-contaminated test set (nc-test) that contains all pairs with patents published in 2024 (n=71), i.e., after the pretraining cut-off date of all evaluated open-weight LLMs, addressing the concern that LLMs might have seen test data during pretraining Ravaut et al.
+(2024).
+3.3 Corpus Statistics and Analysis
+Dataset statistics are presented in Table 1 (further statistics and plots in Appendix K).
+Papers typically include details and analyses regarding experiments; patents usually contain more information on applications and practical benefits.
+We analyze the lexical overlap between the respective documents and find that only 8.3% of the 4-grams are shared.
+This highlights the complexity of the task: the two documents describe the same invention from different perspectives, using different linguistic styles.
+3.4 Proposed Evaluation Metrics
+We adapt a suite of metrics to analyze the performance of patent generation models from multiple perspectives.
+The length and specialized domain of patent documents make automatic evaluation challenging.
+We thus propose to disentangle factual content overlap from stylistic similarity using established metrics that are well-suited for long documents.
+3.4.1 Content-level metrics
+To assess the factual consistency and coverage of the generated patent, we measure semantic overlap with the reference patent and the provided paper.
+The NLI-based metric SCALE Lattimer et al.
+(2023) estimates factual consistency between two documents by computing pairwise entailment scores between premise chunks and hypothesis sentences.
+While premise chunks can be large, SCALE still requires a quadratic computation.
+To make this feasible on long documents, we sample 10 sentences from every hypothesis document.
+For each sentence, we retrieve the 5 most relevant premise chunks according to BM25 and then compute the maximum NLI score across these chunks.
+For system-level scores, we average the scores obtained for all sampled sentences.
+For Pap2Pat, we propose two variants of SCALE:
+Factuality.
+RefGen and Ref+PapGen quantify to what extent the semantic content of the generated patent (Gen) is supported by the reference patent (Ref), and by the reference patent and the paper (Ref+Pap), respectively.
+Coverage.
+To measure the degree to which the generated patent covers the information content of the reference patent, we use the generated patent as premise document and the reference patent as the hypothesis (GenRef).
+We confirm the applicability of these scores to our benchmark by computing a few trivial baselines (see Table 3): taking the reference patent itself as the hypothesis document results in an upper bound of around 89%; using the most similar patent from the train set according to BM25 similarity to the paper results in a score of less than 30%.
+Taken together with the low standard deviations across 5 executions, this demonstrates that the score is able to differentiate and that the sample size is sufficient to make meaningful system-level comparisons.
+Figure 2: Overview of chunk-based outline-guided patent generation (COPGen).
+We chunk the outline, retrieve the parts of the paper that are most relevant for that chunk, prompt the LLM, and concatenate the results.
+The desired output length and the number of allocated patent tokens per chunk determine the number of chunks.
+3.4.2 Language-level metrics
+Patents are written in special style, including the use of partially legally relevant phrases and constructions.
+To estimate to what extent the generated patents adhere to this style, we compare corpus-wide n-gram profiles, an established method for authorship attribution Keselj et al.
+(2003); Zecevic (2011); van Dam (2013); Mikros and Perifanos (2013).
+Here, we use the set of all reference patents from the same split as basis to extract the profile of a typical patent attorney $P^{n}_{\text{ref}}$ and compare it to the profile extracted from the generated patents $P^{n}_{\text{gen}}$.
+Each profile contains the 1000 most frequent n-grams for $n\in\{1,2,3,4\}$.
+To compute the similarity between $P^{n}_{\text{ref}}$ and $P^{n}_{\text{gen}}$, we adopt the formula from Keselj et al.
+(2003):
+$sim_{n}=\sum_{g\in P^{n}_{\text{ref}}\cup P^{n}_{\text{gen}}}1-\left(\frac{P^{n}_{\text{ref}}(g)-P^{n}_{\text{gen}}(g)}{P^{n}_{\text{ref}}(g)+P^{n}_{\text{gen}}(g)}\right)^{2}$
+We additionally integrate a profile from StyloMetrix [^4] https://github.com/ZILiAT-NASK/StyloMetrix Okulska et al.
+(2023), which implements rule-based detection of 196 linguistic features, including e.g.
+verb tenses, modal verbs, POS tags, lexical items, figures of speech, and linguistic constructions such as fronting or similes.
+We compute the final score as the average of the four n-gram profile similarities and the StyloMetrix profile similarity, i.e., $Style=\frac{1}{5}((\sum_{i=1}^{4}{sim_{n}})+sim_{stylo})$.
+To measure the repetitiveness, we compute the repetition rate RR Cettolo et al.
+(2014), i.e., the fraction of n-grams that appear more than once.
+We compute RR over sliding windows of 256 tokens and average the scores.
+To differentiate between repetitive language and unintended infinite repetitions, we additionally report the fraction of windows with an RR score greater than 80 (RR>80).
+4 COPGen: Chunk-based Outline-guided Patent Generation
+Since LLMs cannot yet generate sufficiently long documents in a single call, we instead choose to generate patents in chunks.
+Figure 2 summarizes the approach.
+We chunk the outline, select the most relevant parts of the paper (paper context) depending on the chunks’ outline bullet points, and prompt an LLM to generate the patent text for that chunk.
+The outlines of previous chunks are included for global document context.
+Finally, we concatenate the generated outputs and apply only a lightweight post-processing to remove duplicate headings at chunk boundaries.
+This framework enables generating a long document in parallel, with customized context per chunk and controllable length.
+Tokens Text Sim $\uparrow$ Content-Level Metrics (SCALE) $\uparrow$ Language $\uparrow$ Repetitions
+Coverage Factuality
+BS R-L GenRef RefGen Ref+PapGen Style DS RR RR>80
+Heuristic Baselines / Skylines
+Reference Patent 18.1k (100%) 100 100 88.6 $\pm$ .15 88.5 $\pm$ .18 88.7 $\pm$ .18 100 100 14.4 0.2
+Similar Patent 30.1k (166.0%) 65.8 33.7 29.4 $\pm$ .37 27.6 $\pm$ .32 27.6 $\pm$ .36 75.3 98.3 12.3 0.1
+Outline 1.4k (7.9%) 56.5 10.1 39.7 $\pm$ .72 61.6 $\pm$ .14 61.9 $\pm$ .15 24.4 85.6 20.1 0.2
+Paper 8.1k (44.5%) 69.7 39.4 44.8 $\pm$ .61 46.5 $\pm$ .40 89.0 $\pm$ .13 47.2 98.4 8.4 0.0
+[…]
+Mixtral-8x7B 3.2k (17.7%) 66.1 23.1 38.1 $\pm$ .70 66.9 $\pm$ .87 72.0 $\pm$ .71 42.6 96.9 17.9 0.6
+Qwen2-72B 2.8k (15.6%) 66.6 21.3 40.3 $\pm$ .45 65.8 $\pm$ .47 72.4 $\pm$ .43 39.6 97.3 9.5 0.5
+   w/o Paper 3.5k (19.3%) 66.1 23.2 38.9 $\pm$ .29 65.9 $\pm$ .54 66.6 $\pm$ .38 37.1 96.6 9.8 0.6
+   w/o Outline 2.0k (11.1%) 64.2 16.9 34.9 $\pm$ .40 56.7 $\pm$ .31 75.3 $\pm$ .23 39.8 97.4 9.3 0.1
+[…]
+Llama-3 8B 9.6k (53.1%) 68.7 41.4 40.3 $\pm$ .21 60.8 $\pm$ .49 65.7 $\pm$ .55 43.2 97.0 27.0 4.4
+Llama-3 8B SFT 27.5k (151.5%) 70.4 42.8 42.0 $\pm$ .25 49.3 $\pm$ .39 52.1 $\pm$ .39 59.4 98.0 53.7 29.3
+   w/ rep. removal 17.3k (95.4%) 71.2 49.6 42.1 $\pm$ .48 49.6 $\pm$ .38 53.4 $\pm$ .59 64.7 98.5 38.4 8.5
+Mixtral-8x7B 5.6k (31.0%) 69.1 35.5 41.8 $\pm$ .61 62.3 $\pm$ .31 68.5 $\pm$ .30 49.3 97.5 14.5 0.4
+Llama-3 70B 6.1k (33.9%) 70.2 39.0 42.7 $\pm$ .60 64.5 $\pm$ .47 68.6 $\pm$ .58 49.5 97.4 17.5 0.2
+Qwen2-72B 8.1k (44.8%) 70.2 41.3 44.1 $\pm$ .32 62.5 $\pm$ .44 67.9 $\pm$ .35 47.5 97.3 12.8 1.2
+[…]
+Qwen2-72B 18.1k (100%) 71.7 50.8 46.8 $\pm$ .31 59.7 $\pm$ .63 65.3 $\pm$ .44 47.8 97.1 10.0 0.5
+Table 3: Experimental Results.
+Italic values represent upper bounds that used test data in the prediction.
+The best value per column is bold, the best per section bold italics.
+Tokens are reported as absolute and relative to the reference.
+BS = BERTScore.
+R-L = ROUGE-L.
+DS = DiscoScore.
+Text Sim = Standard Text Similarity Metrics.
+[…]
+[…]
+Paper Context Selection.
+For each chunk $i$, the retriever selects relevant paragraphs from the paper to create the paper context $\text{c}_{i}$.
+We use BM25 Robertson and Zaragoza (2009) with the chunk’s outline $\text{o}_{i}$ as query.
+We always include the abstract of the paper and all headings, adding paper paragraphs successively in the order of their relevance ranking until the token limit is reached.
+Patent Generation.
+The paper context $\text{c}_{i}$, the current outline $\text{o}_{i}$ and prior outlines $\text{o}_{j<i}$ are combined to a prompt (see Appendix J).
+The LLM generates patent chunk $\text{p}_{i}$, using constrained decoding to adhere to the outline’s headings.
+5 Experiments
+In this section, we describe our experiments using Pap2Pat to assess the capabilities and limitations of current LLMs with COPGen.
+5.1 Evaluation Metrics
+We primarily focus on the evaluation metrics proposed in Section 3.4.
+For the sake of completeness, we also report ROUGE-L F1 Lin (2004) and BERTScore F1 Zhang et al.
+(2020).
+For BERTScore, we circumvent the limit of 512 tokens with a sliding window, and use SciBERT Beltagy et al.
+(2019), which has been shown to be effective in the patent domain Pujari et al.
+(2021).
+We also report DiscoScore Zhao et al.
+(2023), a state-of-the-art coherence metric rooted in Centering Theory Grosz et al.
+(1995) and based on BERT (Devlin et al., 2019).
+We use the DS-SENT-NN variant since the authors report that it performs best on long texts.
+We extend the published implementation [^6] https://github.com/AIPHES/DiscoScore with a sliding window to obtain BERT embeddings.
+5.2 Models
+Choice of LLMs.
+To make our work reproducible, we only leverage recently published open-weight LLMs with state-of-the-art results on other generation tasks.
+We include Llama-3 8B and 70B Dubey et al.
+(2024), Mixtral-8x7B Jiang et al.
+(2024), and Qwen2-72B Yang et al.
+(2024).
+Baselines.
+We include four baselines, upper bounds, and sanity checks for the evaluation setup.
+In particular, we consider the paper, the outline, the reference patent and the most similar patent from the train set (highest BM25 similarity to the paper) as the generated document.
+Mixtral-8x7B and Qwen2-72B support 32k tokens, thus we use them to determine the performance without COPGen by prompting them with the complete paper and outline (Single LLM-call).
+Fine-tuning.
+We fine-tune Llama-3 8B on the training split of Pap2Pat using LoRA Hu et al.
+(2022).
+We adopt the hyperparameters proposed by Tribes et al.
+(2024), see Appendix G.
+Since the fine-tuned model frequently generates infinite repetitions, we design a post-processing procedure that detects and removes them (see Appendix C).
+5.3 Main Findings
+Table 3 shows our main evaluation results.
+Overall, we find that COPGen strongly improves upon using a Single LLM-call, in which the generated patents are much too short (on average less than one fifth of the reference patents).
+COPGen provides an effective remedy, creating much longer and length-controllable outputs with higher coverage (correlating with higher text similarity scores).
+In terms of language style, COPGen also has a clear advantage.
+For RR, values around 12-14% seem to be ideal according to the scores of the reference and similar patents.
+While the fine-tuned models produce more repetitions, both the Single LLM-call and zero-shot COPGen have similar repetition rates and only the smaller and fine-tuned Llama models get stuck in infinite repetitions.
+All models achieve very high coherence scores, indicating that all LLMs in the study do not suffer from generating incoherent text.
+[…]
+Length Control and Coverage-Factuality Tradeoff.
+Figure 3 shows that allocating fewer tokens per chunk for the patent, i.e., generating more chunks, leads to a near-linear increase in output length.
+Our length control mechanism is able to find an optimal setting in which the average output length corresponds to that of the reference patents (last row of Table 3), resulting in the overall best text similarity and coverage scores, while keeping factuality high (as opposed to the fine-tuned models that produce similar length).
+Coverage and factuality are intuitively antagonistic (see Figure 3), similar to precision and recall: as the generated text becomes longer, it becomes increasingly difficult to maintain factuality but easier to achieve higher coverage.
+This also explains why the Single LLM-call baselines achieve higher factuality scores than COPGen runs that generate more than five times as much text.
+5.4 Analysis of COPGen Settings
+We now study various settings of COPGen.
+Impact of Outline Granularity.
+In Figure 4, we observe a consistent improvement across coverage, factuality and BERTScore with more detailed outlines: users can directly improve output quality by providing more details.
+Notably, contrary to our experiments on length control, the improvements in coverage are achieved without longer outputs and without decreasing factuality.
+00202020204040404060606060Mean Number of Chunks per Patent555510101010151515152020202025252525Length (k tokens)4242424248484848545454546060606066666666SCALELength (k tokens)CoverageFactuality
+[…]
+Impact of Selection of Paper Context.
+We study how providing informative context taken from the paper influences results.
+In Figure 5, we show retriever ablation results including two baselines: NoPaper does not add any context from the paper, and AbstractOnly uses only the abstract.
+As an upper bound, we use BM25Oracle, where the BM25 query is the original patent text instead of the outline.
+We observe a monotonically increasing performance, demonstrating that associated papers provide valuable information.
+There is only a small gap between BM25 and BM25Oracle, suggesting that the outline is an effective BM25 query and that more elaborate retrieval methods could close the gap further.
+Test Data Contamination.
+If patents are (partially) memorized during pre-training, one would expect a sudden drop in performance when evaluating on patents published after the pre-training cutoff date, i.e., on the non-contaminated test set.
+We see only minor differences in text similarity and SCALE metrics, though the style scores drop significantly (see Appendix D).
+This is likely due to domain distribution shifts (see Appendix K).
+Overall, the results do not indicate systematic issues with memorization.
+Fine-tuning.
+Fine-tuning results in significantly longer outputs, which is consistent with the findings of Bai et al.
+(2024a).
+However, this increased length is partially due to infinite repetitions, an issue also observed in concurrent work on patent generation by Wang et al.
+(2024c).
+Prior work has identified repetitive training data as a key factor contributing to infinite repetitions Li et al.
+(2023a).
+We hence hypothesize that the cause is patents’ inherently repetitive style.
+Patents often present numerous variations of the invention, each introduced with similar phrasing and detailing different combinations of components.
+This also results in a 71% higher RR than for papers.
+Despite these repetitions, fine-tuning improves standard text similarity metrics like ROUGE-L and BERTScore.
+Wang et al.
+(2024c) find the same, along with much lower scores in their human evaluation, concluding that the repetitions lead to over-rewarding in n-gram-based metrics.
+We find contrary evidence that removing many repetitions from the generated patents in post-processing actually further improves ROUGE-L by 6.8pp.
+Our evaluation instead reveals that the improvement achieved by fine-tuning is likely mainly due to stylistic similarity: style scores increase by over 20pp.
+while factuality metrics decrease by over 10pp.
+Coverage4040404045454545Factuality57.557.557.557.560.060.060.060.062.562.562.562.5BERTScore6969696970707070Tokens (k)6666888810101010shortmediumlong
+Figure 4: Outline granularity ablation of Qwen2-72B with default token allocation.
+Coverage40.040.040.040.042.542.542.542.545.045.045.045.0Factuality616161616262626263636363BERTScore6868686870707070ROUGE-L3535353540404040NoPaperAbstractOnlyBM25BM25Oracle
+Figure 5: Retriever ablation. The plot depicts the results of Qwen2-72B using the long outline and default token allocation.
+6 Human Evaluation
+To gain further insight into the practical applicability beyond the automatic metrics, we conduct a human case study with two patent attorneys for the field of AI.
+We randomly select 10 samples from the NLP field out of the Pap2Pat test set and use the outputs of Qwen2-72B with COPGen and default token allocation.
+For each of the samples, the attorneys are provided with the paper, the outline, the generated patent and the reference patent.
+They are then tasked to evaluate the quality of the generated patent based on the hypothetical scenario that they got the paper, wrote the outline and sent both to the LLM.
+We ask about strengths, weaknesses and potential for time savings.
+The attorneys evaluated 10 and 5 samples, respectively, yielding 15 total evaluations and spending about 30 minutes per sample on average.
+Among the 15 evaluations, the attorneys saw substantial time savings in 8 cases, showing promising potential to increase their productivity.
+However, they also identify two main limitations that could hinder such time savings, which are consistent with observations from our automatic evaluation, further validating our chosen metrics.
+These limitations are:
+  1. (1)
+    Style: The model often fails to use non-limiting language, which is essential in the patent description to avoid unnecessary limitations.
+For instance, stating that “the system includes a crucial component” may be too limiting, whereas rephrasing it as “the system may include a preferred component” provides more flexibility to adapt claims in the grant procedure and does not narrow the claims’ interpreted scope of protection.
+Furthermore, the model occasionally uses promotional phrases like “have revolutionized” that are common in NLP papers but inappropriate in patents.
+  2. (2)
+    Level of detail: The patent description should describe the invention precisely enough to enable an ordinary person skilled in the art to reproduce it.
+This level of detail is often still not generated, which is also in line with the lower absolute scores for coverage compared to those for factuality.
+7 Conclusion
+In this work, we have presented the first study on the generation of complete patent descriptions in a realistic experimental setting, using papers as real-world invention specifications.
+We create the Pap2Pat benchmark, propose targeted evaluation metrics, develop COPGen to enable the generation of long patent documents, and conduct a human evaluation.
+We find that LLMs can effectively generate patent drafts from research papers, but that there is still headroom for improving the level of detail and linguistic style.
+Promising directions for future research include developing fine-tuning methods that maintain factuality and avoid repetitions, as well as conducting user studies to identify efficient interaction patterns.
+Limitations
+This study focuses on open-weight models, leaving the exploration of closed-source commercial models as a potential avenue for future research.
+Expanding our experimental setting to include these models could provide additional insights.
+In this work, we do not study the automatic generation of outlines, but consider them user input.
+Generating them in a planning stage from the paper or other input could help decrease manual effort further.
+Due to the immense cost of patent attorneys, our expert case study analyses only a small sample of the dataset. Conducting large-scale human evaluations could generate further insights.
+Acknowledgments
+We would like to thank the patent attorneys Philipp Mangold and Charlotte Hellmann for evaluating the quality of generated patents, and for providing valuable insights into the peculiarities of patents.
+Ethics Statement
+The development of AI-powered patent generation tools may raise ethical concerns.
+We emphasize that our research is intended to support and augment the work of patent professionals, rather than to fully automate the patenting process or facilitate malicious activities such as patent trolling.
+Appendix A Pap2Pat Dataset
 Field Patent Paper Candidate 1 () Paper Candidate 2 ()
@@
 Table 4: Example for the matching of a PPP.
-The scores correspond to the respective metrics, the term metrics are shown as min-normalized / max-normalized.
+The scores correspond to the respective metrics, the term metrics are shown as $\min$-normalized / $\max$-normalized.
 Both papers have author lists that contain all the inventors, were published inside the date range, and have titles and abstracts with sufficiently high absolute term similarity to the patent.
 Since the term similarity scores of paper 1 are higher than those of paper 2 by the specified margin (see Distinctiveness filtering step), paper 1 is correctly matched to the patent.
-3 Pap2Pat Dataset
-In this section, we present the Pap2Pat dataset containing 1.8k PPPs from a variety of domains, each annotated with multiple outlines.
-It serves two main purposes: (1) It is an extremely challenging benchmark for LLMs because it contains very long and complex patent documents, and requires deep understanding of the technical domain as well as patent law.
-(2) It facilitates the development of AI-powered tools for patent drafting, where the patent attorney only performs post-editing rather than writing from scratch, potentially incurring massive cost savings.
+We present the Pap2Pat dataset containing 1.8k PPPs from a variety of domains, each annotated with multiple outlines.
+It serves two main purposes.
+First, it is a challenging benchmark for LLMs that requires long-text generation capabilities and deep understanding of the technical domain as well as patent law.
+Second, it facilitates the development of AI-powered tools for patent drafting, where the patent attorney only performs post-editing rather than writing from scratch, potentially incurring massive cost savings.
 In the following, we describe the steps taken for constructing the Pap2Pat dataset: scraping and filtering PPPs, parsing the full-text documents and generating the patent outlines.
-3.1 Scraping Patent-Paper Pairs
-To create Pap2Pat, we start out with the USPTO dataset [^4] https://bulkdata.uspto.gov containing 6.7M patent applications from 2005 to April 2024.
-For each patent, we query SemOpenAlex (Färber et al.
-2023) using SPARQL and retrieve papers with overlapping authors lists and publication dates.
+A.1 Scraping Patent-Paper Pairs
+The patent and paper in a PPP typically do not cite each other, so we cannot rely on front-page or in-text citations Marx and Fuegi (2020, 2022) to find PPPs.
+The matching must rely on document metadata and content alone.
+We start out with the USPTO dataset [^7] https://bulkdata.uspto.gov containing 6.7M patent applications from 2005 to April 2024.
+For each patent, we query SemOpenAlex Färber et al.
+(2023) using SPARQL and retrieve papers with overlapping authors lists and publication dates.
 Next, we filter the results based on their titles, abstracts, other candidate matches for the same patent, and paper licenses, as elaborated below.
 Table 2 shows the remaining number of candidates after each filtering step.
+As can be seen, the major limiting factor for our benchmark size is the issue of scientific articles being published under restrictive licenses.
 An example match is shown in Table 4.
-We perform a systematic manual post-hoc evaluation of the heuristics to validate their precision (see Section 3.2).
-Filter Candidates
-Authors + Date 930k
-+ Term Overlap 100k
-+ Distinctiveness 21k
-+ Permissive License 1.8k
-Table 2: Filter criteria and number of candidates.
-Split # pairs # patent tokens # paper tokens # outline bullets
-short medium long
-train 1000 17808.9 $\pm$ 15132.5 7991.7 $\pm$ 4493.4 36.8 $\pm$ 29.2 73.5 $\pm$ 60.0 149.0 $\pm$ 122.0
-val 242 18201.3 $\pm$ 16283.5 8018.5 $\pm$ 4089.2 37.4 $\pm$ 30.8 74.4 $\pm$ 62.9 150.6 $\pm$ 127.5
-test 500 18144.9 $\pm$ 13159.5 8081.3 $\pm$ 3919.8 37.5 $\pm$ 24.7 74.9 $\pm$ 50.9 151.6 $\pm$ 103.7
-nc-test 71 18415.2 $\pm$ 13995.4 9520.3 $\pm$ 4648.0 37.8 $\pm$ 22.7 76.0 $\pm$ 46.4 154.4 $\pm$ 94.8
-all 1813 17977.7 $\pm$ 14724.4 8079.8 $\pm$ 4302.0 37.1 $\pm$ 28.0 74.1 $\pm$ 57.5 150.1 $\pm$ 117.0
-Table 3: Dataset Statistics. Values are reported as mean $\pm$ std. Token counts correspond to the Llama-3 tokenizer.
+We perform a systematic manual post-hoc evaluation of the heuristics to validate their precision (see Section A.2).
 Author Overlap.
 The patent and the paper of a PPP are by definition authored by overlapping sets of individuals.
-The overlap of author lists have therefore been identified as an effective (yet not sufficient) criteria for matching PPPs (Magerman, Van Looy, and Song 2010).
-The requirements for paper authorship are typically much lower than those for patent inventorship (Konski and Wu 2015).
+The overlap of author lists have therefore been identified as an effective (yet not sufficient) criteria for matching PPPs Magerman et al.
+(2010).
+The requirements for paper authorship are typically much lower than those for patent inventorship Konski and Wu (2015).
 In many cases, only the main author(s) and the senior author(s) are listed as inventors.
@@
 This score’s effectiveness increases with the number of inventors, so we only consider patents with at least two inventors.
-Implementating $\text{sim}_{\text{author}}$ requires some form of author name disambiguation to avoid false negatives (different spellings, e.g., with, without, or with abbreviated middle name) and false positives (e.g., very common names like John Smith).
+Implementing $\text{sim}_{\text{author}}$ requires some form of author name disambiguation to avoid false negatives (different spellings, e.g., with, without, or with abbreviated middle name) and false positives (e.g., very common names like John Smith).
 There are no author identifiers shared between the patent and paper datasets, so our disambiguation uses the surface names only.
@@
 We compare titles and abstracts of patents and papers using term overlap metrics.
-We first obtain a set of terms $T$ using removal of stopwords and punctuation666based on spaCy’s en_core_web_sm, and stemming777based on NLTK’s PorterStemmer.
-The score is then computed as the number of shared terms, normalized by the minimum or maximum number of terms, following Magerman, Looy, and Debackere (2015).
-We additionally weight each term using its IDF value across all candidate titles and abstracts.
-We compute these IDF weights based on the titles and abstracts of all candidate patents and papers.
-$\text{sim}_{\text{term}}(s_{1},s_{2})$ $=\frac{\displaystyle\sum_{t\in T(s_{1})\cap T(s_{2})}\text{idf}(t)}{\displaystyle\text{agg}\Big{(}\sum_{t\in T(s_{1})}\text{idf}(t),\sum_{t\in T(s_{2})}\text{idf}(t)\Big{)}}$
-where $s_{1}$ and $s_{2}$ are either the titles or the abstracts of the candidate pair, and $\text{agg}\in\{\min,\max\}$.
+We first obtain a set of terms $T$ using removal of stopwords and punctuation999based on spaCy’s en_core_web_sm and stemming.101010based on NLTK’s PorterStemmer The score is then computed as the number of shared terms, normalized by the minimum or maximum number of terms, following Magerman et al.
+(2015).
+We additionally weight each term using its IDF value across all titles and abstracts of patents and papers.
+$\text{sim}_{\text{term}}(pat,pap)$ $=\frac{\displaystyle\sum_{t\in T(pat)\cap T(pap)}\text{idf}(t)}{\displaystyle\text{agg}\Big{(}\sum_{t\in T(pat)}\text{idf}(t),\sum_{t\in T(pap)}\text{idf}(t)\Big{)}}$
+where $pat$ and $pap$ are either the titles or the abstracts of the documents, and $\text{agg}\in\{\min,\max\}$.
 Thus, we have 4 scores in total, for which we set the thresholds to be 0.15 with $\min$ normalization and 0.1 with $\max$ normalization.
 We choose the values based on interactive experimentation, but perform a post-hoc validation of the matching precision.
-Figure 2: Our approach to chunk-based outline-guided generation. The desired section length information is used to determine chunk sizes for creating outline chunks $o_{1}$, $o_{2}$, …, $o_{n}$.
 Distinctiveness.
@@
 License.
-We filter the matches for licenses that allow redistribution and commercial use, i.e., CC-BY, [^8] https://creativecommons.org/licenses/by/4.0/ CC0, [^9] https://creativecommons.org/publicdomain/zero/1.0/ and public domain.
+We filter the matches for licenses that allow redistribution and commercial use, i.e., CC-BY,CC0,and public domain.
 We use the license information provided by SemOpenAlex and by the ArXiv API.
-3.2 Manual Validation
+A.2 Manual Validation
 To verify the precision of our matching pipeline, we perform a manual validation, conducted by the first author of this paper.
@@
 In the five imperfect matching cases, the papers still provide meaningful training and evaluation signals, as they are still closely related to the invention and contextualized by the outline.
-3.3 Document Parsing
+A.3 Document Parsing
 We parse all patents and papers into a nested JSON schema where each section has a title, paragraphs and subsections field.
-We make considerable efforts to obtain clean data: we perform LLM-based section hierarchy reconstruction for patents, font-based section hierarchy reconstruction for paper PDFs and formula conversion for patents and paper XMLs from PubMed.
-We provide more details in Appendix G.
-3.4 Outline Generation
-One key aspect of our approach are the patent outlines, which simulate the patent attorney’s input to the system in production settings.
-Examples are shown in Figure 2 and Appendix A.
-They provide the model with the target headings, as well as bullet points summarizing the document structure and high-level content of every section.
-For automatically generating the outlines for Pap2Pat, we leverage the original patents and Llama-3 70B (Dubey et al.
-2024).
-We provide the prompt in Appendix E.
-To ensure that the model adheres to the output format, we use SGLang (Zheng et al.
-2023) for constrained decoding.
-We enforce a fixed number of bullet points $n_{\text{bullets}}$ per section, where $n_{\text{bullets}}$ is proportional to the length of the text in that section ($n_{\text{chars}}$).
-$n_{\text{bullets}}=\begin{cases}\text{max}(1,\lfloor n_{\text{chars}}/l\rfloor)&\text{if }n_{\text{chars}}>0\\ 0&\text{else}\end{cases}$
-where $l$ is the number of characters that each bullet point summarizes on average.
-We create outlines in three levels of granularity: long ($l$=500, avg.
-150 items), medium ($l$=1000, avg.
-74 items) and short ($l$=2000, avg.
-37 items), see Table 3.
-For each section (identified by a heading), we additionally provide the number of characters in the original patent, which can function as the desired content lengths during generation.
-In a practical application setting, the patent attorney would provide them alongside the outlines.
-3.5 Dataset Splits and Statistics
+We make considerable efforts to obtain clean data: we perform LLM-based section hierarchy reconstruction for patents, font-based section hierarchy reconstruction for paper PDFs and formula conversion for patents and papers.
+We provide more details in Appendix B.
+A.4 Dataset Characteristics
 We split our dataset randomly into train (n=1000), test (n=500) and validation (n=242).
 We additionally create a non-contaminated test set (nc-test) that contains all pairs with a patent published in 2024 (n=71), i.e., after the pretraining cut-off date of all evaluated open-weight LLMs.
-Thus, we address the concern that LLMs might have seen test data during pretraining (Ravaut et al.
-2024).
-Table 3 shows dataset statistics across the splits, further statistics and plots are presented in Appendix H.
-Dataset Split test nc-test
-Metric BERTScore ROUGE-L Tokens BERTScore ROUGE-L
-Outline short medium long short medium long long long long
-Llama-3 8B 67.8 68.3 68.7 23.7 23.0 23.9 51.3% 68.8 22.7
-Llama-3 70B 68.6 69.5 70.2 24.7 23.2 24.8 33.6% 70.5 24.1
-Mixtral-8x7B 68.1 68.3 69.1 23.6 23.1 25.4 31.7% 68.8 24.3
-Qwen2-72B 69.0 69.4 70.2 25.7 24.9 26.1 42.9% 70.4 25.1
-Llama-3 8B SFT 69.2 69.9 70.5 26.0 27.2 26.5 162.2% 71.6 26.9
-Table 4: Experimental Results. SFT stands for supervised fine-tuned. For the number of tokens and metrics on nc-test, we only show the results regarding the long outline due to space constraints.
-4 Chunk-based Outline-guided Generation
-In this section, we propose chunk-based outline-guided generation, a novel approach to patent generation illustrated in Figure 2.
-It serves as a strong baseline for comparison in future work.
-It is motivated by current LLMs’ limitations regarding sequence length (which needs to cover the instruction, the context, and the output) and therefore generates the patent in chunks, guided by the user-defined outline.
-We chunk the outline, retrieve paper context based on the outline chunk, and prompt an LLM to generate the patent text for that chunk.
-Finally, we concatenate the generated outputs and apply only a lightweight post-processing to remove duplicate headings at chunk boundaries.
-Chunking and Token Allocation.
-We assign a fixed number of tokens for the instruction (including system prompt, user prompt and outlines) and equally split the remaining tokens among paper context and output patent.
-We then chunk the outline using the desired length of the respective patent section.
-The chunking procedure is designed to pack as much content as possible into each chunk, while preserving the integrity of individual sections.
-Therefore, we only split a section if it exceeds the token limit on its own.
-Retriever.
-For each chunk $i$, the outline $\text{o}_{i}$ is passed to a retriever, which selects relevant paragraphs from the paper to create the paper context $\text{c}_{i}$.
-We use BM25 (Robertson and Zaragoza 2009) as the retrieval method and the chunk’s outline as the query.
-We always include the abstract of the paper because it provides a valuable overview, and successively add paragraphs in the order of their relevance ranking until the token limit is reached.
-The results are formatted in a Markdown document that comprises all headings and the retrieved paragraphs.
-LLMs.
-The paper context $\text{c}_{i}$, the current outline $\text{o}_{i}$ and prior outlines $\text{o}_{j<i}$ are combined to form the prompt to the LLM (see Appendix F).
-Based on this prompt, the LLM generates the patent chunk $\text{p}_{i}$, where constrained decoding ensures that the model adheres strictly to the headings in the outline.
-To make our work reproducible, we only leverage recently published open-weights LLMs that have been reported to achieve state-of-the-art results on other generation tasks.
-We include a small model (Llama-3 8B (Dubey et al.
-2024)) and several larger models (Mixtral-8x7B (Jiang et al.
-2024), Llama-3 70B (Dubey et al.
-2024) and Qwen2-72B (Yang et al.
-2024)).
-Per default, we use a maximum sequence length of 8k tokens for all LLMs to ensure that the results are comparable and leave the study of the effect of increased maximum sequence length to future work.
-Fine-tuning.
-We fine-tune Llama-3 8B on the training split of Pap2Pat using LoRA (Hu et al.
-2022).
-We use the same prompt and BM25 retrieval as during inference and adopt the hyperparameters proposed by Tribes et al.
-(2024), see Appendix C.
-5 Experiments
-We perform a series of experiments using the Pap2Pat dataset to assess the performance of current open-weight LLMs on the proposed task.
-5.1 Experimental Setup
-To compare the generated and original patents, we use the n-gram-based metric ROUGE-L (F1) (Lin 2004), as well as the neural metric BERTScore (F1) (Zhang et al.
-2020) for a balanced evaluation.
-For BERTScore, we use SciBERT (Beltagy, Lo, and Cohan 2019) as the base model, as it has been shown to be effective in the patent domain (Pujari, Friedrich, and Strötgen 2021).
-To address the context length limitations of BERT architectures, we process the input in overlapping chunks, where each chunk includes a portion of the preceding chunk to provide contextualization across chunk boundaries.
-We then compute the dot products between these embeddings across the whole sequence, as in the original variant.
-The column titled Tokens reports the ratio of the number of generated tokens and number of tokens of the original patent using the Llama-3 tokenizer.
-5.2 Results
-Table 4 shows the main evaluation results.
-The three larger models all outperform Llama-3 8B.
-However, all models generate patents that are much shorter than the original ones, even though the prompt includes the desired length.
-BERTScore686868687070707072727272ROUGE-L2323232324242424252525252626262627272727Tokens353535354040404045454545NoPaperAbstractOnlyBM25BM25Oracle
-Figure 3: Ablation of retrievers. The plot depicts the results of Qwen2-72B using the long outline.
-CS (n=130) Bio (n=152)
-BS RL BS RL
-Llama-3 8B 67.5 22.1 69.4 25.1
-Llama-3 70B 69.0 23.1 70.9 26.2
-Mixtral-8x7B 67.5 23.5 70.2 26.7
-Qwen2-72B 69.0 24.8 71.0 27.5
-Llama-3 8B SFT 70.2 26.2 71.0 26.7
-Table 5: Performance comparison between domains (CS=Computer Science, Bio=Biology). The metrics are BERTScore (BS) and ROUGE-L (RL).
-Outline Granularity.
-With all models, the long outline achieves the highest BERTScore and ROUGE-L.
-We observe a monotonically increasing BERTScore with more detailed outlines.
-The same is not true for ROUGE-L, where the short outline consistently achieves higher scores than the medium one.
-Overall, the performance differences between outlines of different granularity are rather small, showcasing the robustness and flexibility of the proposed approach.
-Fine-tuning.
-BERTScore and ROUGE-L improve substantially through fine-tuning, even beyond the scores of the larger models.
-It is also notable that the generated patents are more than 3 times longer than before fine-tuning.
-However, the increased length is at least partially due to increased proneness for generating repetitions, which we further discuss in Section 6.
-Retrieval.
-In Figure 3, we show ablation results for the retriever.
-We include two baselines: NoPaper does not add any context from the paper, and AbstractOnly uses only the abstract.
-As an upper bound, we use BM25Oracle, where the BM25 query is the original patent text.
-We find that adding paper context substantially improves performance across metrics, demonstrating that associated papers provide valuable information.
-We observe a monotonically increasing performance across models and outline granularities.
-There is a small yet consistent gap between BM25 and BM25Oracle.
-This suggests that using the outline as a BM25 query is an effective approach, and that more elaborate retrieval methods could close that gap further.
-Differences between Domains.
-We analyze the performance across domains and show the results in Table 5.
-We include the two most represented domains in the dataset: computer science and biology.
-All models perform slightly better on patents from biology, with the difference being more pronounced in ROUGE-L than in BERTScore.
-Test Data Contamination.
-We use the non-contaminated test set (nc-test) to study the effect of potential pre-trained memorization on the task performance.
-If patents are (partially) memorized, one would expect a sudden drop in performance when evaluating on patents published after the pre-training cutoff date.
-We observe an equal or slightly increased BERTScore and a slightly decreased ROUGE-L, which are more likely attributed to random noise or domain shifts (see Appendix H), and do not indicate systematic issues with LLM memorization affecting our results.
-6 Analysis
-In this section, we provide deeper insights into the relationship between patents and papers and common error patterns.
+Thus, we address the concern that LLMs might have seen test data during pretraining Ravaut et al.
+(2024).
+Table 1 shows dataset statistics across the splits.
+Appendix K shows further statistics and plots, including the distribution over domains and the number of pairs over time.
 In general, both patents and papers contain information not present in the other.
@@
 Nevertheless, we find that it is common for attorneys to copy content from the paper to the patent (or vice versa).
-For instance, many patents and papers share a portion of the figures.
-But text is also commonly copied between the documents.
-We find that the longest common substring between paper and patent contains 27 words on average, with a maximum of 495 words.
-Model-generated patents and corresponding papers, on the other hand, only have a longest common substring of 15 words on average.
-We furthermore observe that if the paper gives the proposed method or invention a name, the model tends to use that name in the patent as well, whereas the original patents often do not.
-The most prevalent error pattern in the generation process is repetition.
-The models sometimes get stuck in a loop where they repeat the same sentence or paragraph indefinitely.
-This issue is most pronounced with Llama-3 8B, in particular after fine-tuning, but also occurs with the larger models.
-We quantify repetitions using RR (Cettolo, Bertoldi, and Federico 2014), a metric that measures the fraction of n-grams that appear just once in a text (lower is better).
-The patents generated by Llama-3 8B have a much higher RR (53.2 before fine-tuning and 58.9 after fine-tuning), compared to the original patents (42.9) and larger models (Llama-3 70B: 44.1, Qwen2-72B: 40.6, Mixtral-8x7B: 42.3).
-We hypothesize that this issue is rooted in the inherent repetitiveness of patents.
-It is common that every paragraph starts with the same words for pages (e.g., ”In some embodiments, component A comprises …”).
-This is also reflected in the fact that patents have a much higher RR (42.9) than papers (34.5) on our test set.
-When generating such repetitive text, the model is more likely to get stuck in a loop of repetitions (Holtzman et al.
-2020; Fu et al.
-2021; Li et al.
-2023a).
-We attempt to alleviate the issue by increasing the temperature and including a repetition penalty.
-We find that these approaches decrease the number of repetitions but also compromise task performance.
-Developing specialized methods to avoid exact-repetitions for texts that contain many near-repetitions presents an interesting direction for future research.
-In our experiments, we use 8k as chunk size for all models.
-We also perform preliminary experiments to test whether using a larger chunk size could lead to better and more coherent patents, but find that the repetition problem overshadows these effects.
-For larger chunk sizes, the repetitions remain uninterrupted for longer, meaning that the content of following outline bullet points does not get addressed.
-We conclude that the repetition problem must be mitigated before a longer context window can achieve any benefits in automated patent drafting.
-7 Conclusion
-In this work, we tackle the task of patent generation.
-Our work is the first to study the generation of patent descriptions and to use PPPs to ground the generation process.
-We build the Pap2Pat benchmark dataset for training and evaluation, and propose a chunk-based approach to outline-guided paper-to-patent generation and evaluate it in both zero-shot and fine-tuning settings.
-We find that associated papers provide relevant information, which can be successfully injected into the LLM’s context via a simple retrieval model based on the outline.
-Furthermore, the approach is robust across different outline granularities and domains.
-We identify repetitions as the most prevalent error pattern and hypothesize that it is related to the inherent repetitiveness of patents.
-Promising directions for future research include addressing the repetitions, effectively leveraging models with longer context windows, and utilizing agent frameworks in the generation process.
-Limitations
-This study focuses on open-weight models, leaving the exploration of closed-source commercial models as a potential avenue for future research.
-Expanding our experimental setting to include these models could provide additional insights.
-Another limitation of our work is that we have only utilized text overlap metrics, and not utilized more advanced evaluation methods, like with LLMs as judge, or via a user study.
-While designed to facilitate practical user interaction, the effectiveness of this approach in real-world scenarios remains to be investigated.
-Future studies will be necessary to fully understand the performance and user experience, and identify areas for improvement.
-Ethics Statement
-The development of AI-powered patent generation tools may raise ethical concerns.
-We emphasize that our research is intended to support and augment the work of patent professionals, rather than to automate the patenting process or facilitate malicious activities such as patent trolling.
-Appendix A Patent Outline Example
+For instance, many patents and papers share a portion of the figures, as well as verbatim copied text.
+Appendix B Document Parsing Pipeline
+In total, we use three different data sources for the full texts.
+For the patents, we use the USPTO bulk downloads [^11] https://bulkdata.uspto.gov/.
+For the papers, we use PubMed if available and PDF otherwise.
+The goal is to extract a clean representation of the full text into a common JSON format.
+In this format, every section has a title, a list of paragraphs and a list of subsections.
+We write parsers for the XML formats from USPTO, PubMed and the PDF parser Grobid [^12] https://github.com/kermitt2/grobid.
+In addition, we perform several cleaning steps:
+  1. 1.
+    PDF Hierarchy Reconstruction.
+The JSON format is hierarchical by nature, for instance to enable better chunking of patents and section-based retrieval from papers.
+However, Grobid does not detect section levels if the sections are not numbered.
+To reconstruct the levels in these cases, we implement a solution that searches for the headings in the PDF file, extracts their font properties, orders them by size, boldness and capitalization, and infers the levels from that.
+  2. 2.
+    Patent Hierarchy Reconstruction.
+In USPTO’s patent XML files, there is a level attribute associated with every heading, but we find that it is rarely correct; most headings are placed on the same level.
+To reconstruct the levels, we pass the list of headings to an LLM and instruct it to infer the levels based on their names (e.g., "Example 1" and "Example 2" and likely children of "EXAMPLES").
+  3. 3.
+    Formula Conversion.
+Many patents and papers include formulas that can be an important part of the document.
+However, in USPTO’s and PubMed’s XML formats, formulas are represented in MathML syntax, which is extremely hard to read and arguably hard to generate.
+To that end, we convert all MathML formulas to latex using pandoc [^13] https://github.com/jgm/pandoc.
+  4. 4.
+    Metadata Section Filtering.
+Patents usually contain a number of metadata sections in the full text, such as information regarding funding or cross-references to related patents.
+To filter out these sections, we collect a list of such heading names and remove a section if its heading has a Levenshtein distance less than 3 to any one of the blacklisted headings.
+Appendix C Repetition Removal
+Algorithm 1 Repetition Detection
+1:function matches(l1, l2)
+2:     $n\_match\leftarrow 0$
+3:     for $i\leftarrow 0$ to $\text{len}(l1)-1$ do
+4:         if $l1[i]=l2[i$then
+5:              $n\_match\leftarrow n\_match+1$
+6:         end if
+7:     end for
+8:     return n_match / len(l1) $>$ 0.9
+9:end function
+10:
+11:function detect_repetitions(words, min_length, max_cycle_length)
+12:     $n\leftarrow len(words)$
+13:     for $k=1$ to max_cycle_length do
+14:         if matches(words[-k:], words[-2 * k : -k]) then $\triangleright$ Check for repetition of length k
+15:              $i\leftarrow 2$triangleright Number of times the pattern is repeated
+16:              while matches(words[-k:], words[-(i + 1) * k : -i * k]) do $\triangleright$ Search for additional occurrences
+17:                  $i\leftarrow i+1$
+18:              end while
+19:              remove_indices $\leftarrow$ [n - (i - 1) * k, …, n]
+20:              total_length $\leftarrow i*k$
+21:              if total_length $\geq$ min_length then
+22:                  words $\leftarrow$ words[: n - (i - 1) * k] $\triangleright$ Keep only first occurrence of pattern
+23:                  remove_rest $\leftarrow$ detect_repetitions(words, min_length, max_cycle_length) $\triangleright$ Recursive call for remaining text
+24:                  return remove_indices + remove_rest $\triangleright$ Return indices to remove
+25:              end if
+26:         end if
+27:     end for
+28:     return []
+29:end function
+We design a procedure to remove infinite repetitions from generated outputs to study their effect on evaluation metrics.
+In that context, we characterize an infinite repetition as a sequence of tokens that appears multiple times until the end of the generation.
+To find such repetitions, we apply the following recursive algorithm in Algorithm 1 to each chunk and remove the returned indices.
+To account for reptitions where the model alters the patterns slightly (e.g., incrementing a number) in each iteration, we consider two word sequences equal if 90% of their positions are equal.
+By default, we use min_length = 50 and max_cycle_length = 300.
+Appendix D Contamination Results
+Tokens Text Sim $\uparrow$ Content-Level Metrics (SCALE) $\uparrow$ Language $\uparrow$ Repetitions
+Coverage Factuality
+BS R-L GenRef RefGen Ref + PapGen Style DiscoScore RR RR>80
+Heuristic Baselines / Skylines
+Reference Patent 18.4k (100.0%) 100.0 100.0 88.6 $\pm$ .57 88.5 $\pm$ .23 88.7 $\pm$ .21 99.8 100.0 15.2 0.1
+Similar Patent 26.1k (141.8%) 66.0 35.3 31.4 $\pm$ .93 28.0 $\pm$ .64 28.3 $\pm$ .68 53.9 98.1 12.5 0.1
+Outline 1.4k (7.6%) 56.7 10.2 42.1 $\pm$ .94 63.4 $\pm$ .88 63.8 $\pm$ .86 25.9 85.1 19.4 0.0
+Paper 9.5k (51.7%) 69.6 42.2 46.7 $\pm$ 1.01 46.5 $\pm$ .87 88.8 $\pm$ .41 37.2 98.0 8.2 0.0
+[…]
+Mixtral-8x7B 3.0k (16.3%) 66.5 21.7 40.7 $\pm$ 1.16 67.2 $\pm$ 1.19 73.9 $\pm$ 1.16 36.5 96.8 15.9 0.1
+Qwen2-72B 2.9k (15.5%) 66.4 21.2 42.0 $\pm$ 1.21 65.0 $\pm$ .76 71.9 $\pm$ .23 34.5 96.9 8.9 0.0
+   w/o Paper 3.4k (18.2%) 65.9 21.8 41.4 $\pm$ 1.25 66.3 $\pm$ 1.04 67.0 $\pm$ .99 34.1 96.1 9.0 0.2
+   w/o Outline 2.0k (11.0%) 63.7 15.8 36.7 $\pm$ 1.27 56.5 $\pm$ .67 75.7 $\pm$ 1.13 30.3 96.8 8.2 0.0
+[…]
+Llama-3 8B 9.6k (52.1%) 68.8 41.4 42.3 $\pm$ 1.11 60.3 $\pm$ 1.60 65.6 $\pm$ 1.23 36.8 96.6 25.4 2.7
+Llama-3 8B SFT 27.0k (146.8%) 71.2 44.0 44.0 $\pm$ 1.47 49.4 $\pm$ 1.42 52.1 $\pm$ 1.36 45.0 97.8 53.7 27.6
+   w/ rep. removal 18.2k (99.1%) 72.1 51.1 44.7 $\pm$ 1.03 52.0 $\pm$ 1.13 55.4 $\pm$ 1.15 50.7 98.2 39.6 8.3
+Mixtral-8x7B 6.3k (34.4%) 68.8 34.4 45.1 $\pm$ 2.00 64.4 $\pm$ 1.40 70.6 $\pm$ 1.24 39.5 96.9 14.4 0.1
+Llama-3 70B 6.1k (33.4%) 70.5 39.5 45.9 $\pm$ 1.15 64.7 $\pm$ 1.06 68.7 $\pm$ .87 44.6 97.0 17.2 0.1
+Qwen2-72B 8.2k (44.7%) 70.4 40.1 46.9 $\pm$ 1.26 65.3 $\pm$ .83 69.8 $\pm$ .42 44.3 97.0 10.9 0.0
+[…]
+Qwen2-72B 17.2k (93.2%) 71.5 50.5 49.9 $\pm$ 1.70 59.2 $\pm$ 1.00 64.9 $\pm$ .43 41.9 96.6 8.8 0.2
+Table 5: Experimental Results on the non-contaminated test set.
+Italic values represent upper bounds that used test data in the prediction.
+The best value per column is bold, the best per section bold italics.
+Tokens are reported as absolute and relative to the reference.
+BS = BERTScore.
+R-L = ROUGE-L.
+Text Sim = Standard Text Similarity Metrics.
+Appendix E Patent Outline Example
 1# DESCRIPTION
@@
 Listing 1: Example patent outline (short variant) for the pair W6364285-US20140112582. The outline corresponds to more than 5 pages. This example was randomly selected.
-Appendix B Patent Structure
+Appendix F Patent Structure
 Figure 6: Illustration of the structure of a patent from Jiang and Goetz (2024).
@@
 In our experiments, we exclude sections containing only metadata, such as statements regarding funding.
-Appendix C Hyperparameters
+Appendix G Hyperparameters
 Generation Training
@@
 We estimate the total number of GPU-hours to be 720.
-Appendix D SPARQL Query
+Appendix H SPARQL Query
 1PREFIX fabio: <http://purl.org/spar/fabio/>
@@
 Listing 2: SPARQL query template used to retrieve papers for a given patent from SemOpenAlex. Template variables <var> are filled based on the query patent.
-Appendix E Summary Generation Prompt
+Appendix I Summary Generation Prompt
 1<|begin_of_text|><|start_header_id|>system<|end_header_id|>
@@
 Listing 3: Prompt used to generate bullet point summaries with Llama-3 70B
-Appendix F Patent Generation Prompt
+Appendix J Patent Generation Prompt
 1### ROLE
@@
 Listing 8: User prompt used for outline-guided paper-to-patent generation (Part 4)
-Appendix G Document Parsing Pipeline
-In total, we use three different data sources for the full texts.
-For the patents, we use the USPTO bulk downloads [^10] https://bulkdata.uspto.gov/.
-For the papers, we use PubMed if available and PDF otherwise.
-The goal is to extract a clean representation of the full text into a common JSON format.
-In this format, every section has a title, a list of paragraphs and a list of subsections.
-We write parsers for the XML formats from USPTO, PubMed and the PDF parser Grobid [^11] https://github.com/kermitt2/grobid.
-In addition, we perform several cleaning steps:
-  1. 1.
-    PDF Hierarchy Reconstruction.
-The JSON format is hierarchical by nature, for instance to enable better chunking of patents and section-based retrieval from papers.
-However, Grobid does not detect section levels if the sections are not numbered.
-To reconstruct the levels in these cases, we implement a solution that searches for the headings in the PDF file, extracts their font properties, orders them by size, boldness and capitalization, and infers the levels from that.
-  2. 2.
-    Patent Hierarchy Reconstruction.
-In USPTO’s patent XML files, there is a level attribute associated with every heading, but we find that it is rarely correct; most headings are placed on the same level.
-To reconstruct the levels, we pass the list of headings to an LLM and instruct it to infer the levels based on their names (e.g., "Example 1" and "Example 2" and likely children of "EXAMPLES").
-  3. 3.
-    Formula Conversion.
-Many patents and papers include formulas that can be an important part of the document.
-However, in USPTO’s and PubMed’s XML formats, formulas are represented in MathML syntax, which is extremely hard to read and arguably hard to generate.
-To that end, we convert all MathML formulas to latex using pandoc [^12] https://github.com/jgm/pandoc.
-  4. 4.
-    Metadata Section Filtering.
-Patents usually contain a number of metadata sections in the full text, such as information regarding funding or cross-references to related patents.
-To filter out these sections, we collect a list of such heading names and remove a section if its heading has a Levenshtein distance less than 3 to any one of the blacklisted headings.
-Appendix H Dataset Statistics
+Appendix K Dataset Statistics
 0.00.00.00.00.20.20.20.20.40.40.40.4train0.00.00.00.00.20.20.20.20.40.40.40.4val0.00.00.00.00.20.20.20.20.40.40.40.4test0.00.00.00.00.20.20.20.20.40.40.40.4nc-testBiologyComputer scienceMaterials scienceChemistryMedicinePhysicsEnvironmental scienceEngineeringothers
@@
 Figure 10: Date offsets between patents and papers. Negative offset means paper was published first.
+Style Length Coverage Factuality
+Biology (n=152)
+Llama-3 8B 41.7 12.2k / 23.0k 39.3 $\pm$ 1.4 61.0 $\pm$ 0.5
+Llama-3 70B 48.9 17.7k / 23.0k 42.0 $\pm$ 0.9 66.2 $\pm$ 1.2
+Mixtral-8x7B 47.7 17.4k / 23.0k 41.3 $\pm$ 1.1 63.7 $\pm$ 0.6
+Qwen2-72B 48.3 10.0k / 23.0k 42.9 $\pm$ 1.1 63.2 $\pm$ 1.0
+Computer Science (n=130)
+Llama-3 8B 37.5 17.0k / 13.1k 42.5 $\pm$ 1.2 61.0 $\pm$ 0.6
+Llama-3 70B 44.5 14.6k / 13.1k 44.2 $\pm$ 0.8 63.4 $\pm$ 0.7
+Mixtral-8x7B 44.7 14.0k / 13.1k 43.5 $\pm$ 0.9 61.2 $\pm$ 0.5
+Qwen2-72B 45.4 15.3k / 13.1k 46.7 $\pm$ 0.8 61.6 $\pm$ 1.0
+Table 7: Domain comparison. We report the style score, the length of the generated and reference patents (Length, generated/reference), coverage Gen → Ref, and factuality Ref → Gen.
+Differences between Domains.
+We analyze the performance across domains (see Appendix K for domain distributions) and show the results in Table 7.
+We include the two most represented domains in the dataset: computer science (CS) and biology (Bio).
+Reference patents from the biology domain are much longer than computer science patents.
+We find that generated Bio patents achieve better factuality but lower coverage across models, despite relative lengths being very similar.
+Stylistic similarity is also higher for Bio patents.
+Furthermore, model ranking differ between the domains: while Llama-3 70B performs best on Bio patents, Qwen2-72B is highly competitive on CS patents.
+However, further analysis is needed due to limited sample sizes.